### v3

# Entropy Functions on Two-Dimensional Faces of Polymatroidal Region of Degree Four: Part I: Problem Formulation and More

Abstract
Characterization of entropy functions is of fundamental importance in information theory. By imposing constraints on their Shannon outer bound, i.e., the polymatroidal region, one obtains the faces of the region and entropy functions on them with special structures. In this series of two papers, we characterize entropy functions on the $2$-dimensional faces of the polymatroidal region of degree $4$. In Part I, we formulate the problem, enumerate all $59$ types of $2$-dimensional faces of the region by an algorithm, and fully characterize entropy functions on $49$ types of them. The entropy functions on the remaining $10$ types of faces will be characterized in Part II, among which $8$ types are fully characterized, and $2$ types are partially characterized.
Index terms - entropy function, entropy region, polymatroid, poymatroidal region
I Introduction
Let $N_{n}=\{1,2,\ldots,n\}$ and $\mathbf{X}\triangleq(X_{i},i\in N_{n})$ be a random vector indexed by $N_{n}$. The set function $\mathbf{h}:2^{N_{n}}\to\mathbb{R}$ defined by
$\mathbf{h}(A)=H(X_{A}),\quad A\subseteq N_{n}$
is called the entropy function of $\mathbf{X}$, while $\mathbf{X}$ is called a characterizing random vector of $\mathbf{h}$. The Euclidean space $\mathcal{H}_{n}\triangleq\mathbb{R}^{2^{N_{n}}}$ where entropy functions live is called the entropy space of degree $n$. The set of all entropy functions, denoted by $\Gamma^{*}_{n}$, is called the entropy region of degree $n$. The characterization of entropy functions, i.e., determining whether an $\mathbf{h}\in\mathcal{H}_{n}$ is in $\Gamma^{*}_{n}$, is of fundamental importance in information theory.
Entropy functions are (the rank functions of) polymatroids, i.e., they satisfy polymatroidal axioms, that is, for all $A,B\subseteq N_{n}$,
$\mathbf{h}(A)$ $\geq 0,$ (I.1)
$\mathbf{h}(A)$ $\leq\mathbf{h}(B),\quad\text{ if }A\subseteq B,$ (I.2)
$\mathbf{h}(A)+\mathbf{h}(B)$ $\geq\mathbf{h}(A\cap B)+\mathbf{h}(A\cup B).$ (I.3)
The region in $\mathcal{H}_{n}$ bounded by such inequalities, denoted by $\Gamma_{n}$, is called the polymatroidal region of degree $n$. Thus, $\Gamma_{n}$ is an outer bound on $\Gamma^{*}_{n}$. For more about entropy functions, we referred the readers to [2, Chapter 13-15].
Traditionally, entropy functions are characterized by information inequalities. Those inequalities derived by polymatroidal axioms are called Shannon-type, as they correspond to the non-negativity of Shannon information measures. Since 1998, a series of non-Shannon-type information inequalities, among which Zhang-Yeung inequality is the first one [3], were discovered [4, 5, 6]. Thus $\overline{\Gamma^{*}_{n}}$, the closure of $\Gamma^{*}_{n}$, is strictly included in $\Gamma_{n}$ when $n\geq 4$. Each information inequality determines an outer bound on $\Gamma^{*}_{n}$, as those set functions in $\mathcal{H}_{n}$ dissatisfy it must be located outside $\Gamma^{*}_{n}$. In this series of two papers, we develop a system of entropy function characterization from the perspective of faces of $\Gamma_{n}$, which covers the traditional inequality characterization.
By definition, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$. Thus, each Shannon-type information inequality determines a face $F$ of $\Gamma_{n}$. It is natural to characterize entropy functions on the specific $F$ of $\Gamma_{n}$ (See Subsection II-B for details on the faces of a polyhedral cone). Let $F^{*}\triangleq F\cap\Gamma^{*}_{n}$ be the set of all entropy functions in $F$. In what follows, when we refer to determining the entropy functions on $F$ , or the region $F^{*}$, we will simply call this characterizing $F$. A non-Shannon-type information inequality can be considered as an outer bound on $F^{*}$ when $F=\Gamma_{n}$ itself, the improper face of $\Gamma_{n}$. A constrained non-Shannon-type information inequality is an outer bound on $F^{*}$ when $F$ is the face determined by the constraints that are equalities obtained by setting some Shannon-type inequalities as equalities. When $F$ is an extreme ray, i.e., a $1$-dimensional face of $\Gamma_{n}$, if it contains a matroid, entropy functions on $F$ are called matroidal entropy functions, and they can be fully characterized by the probabilistically characteristic set of the matroid [7, 8, 9]. Matúš fully characterized the first non-trivial 2-dimensional face of $\Gamma_{n}$ in 2006 [10]. It is a $2$-dimensional face of $\Gamma_{3}$. In 2012, Chen and Yeung characterized another $2$-dimensional face of $\Gamma_{3}$ [11]. They are the only two types of non-trivial $2$-dimensional faces of $\Gamma_{3}$ that need to be characterized. To the best of our knowledge, so far, there is no fully characterized no-trivial $3$-dimensional faces. However, partial characterizations of $3$-dimensional faces of $\Gamma_{3}$ can be found in [12, 13].
Many information-theoretic problems can be considered as applications of entropy function characterizations on faces of $\Gamma_{n}$. In a series of three papers [14, 15, 16], Matúš and Studneý solved the probabilistic conditional independence problem for four random variables. Note each class of conditional independence constraints, which is called a semimatroid in their papers, determines a face $F$ of $\Gamma_{n}$. The solution to this problem, i.e., the probabilistic-representability of a semimatroid determines whether the relative interior of the corresponding face $F$ intersets with $\Gamma^{*}_{n}$. This problem thus can be considered as partial characterizations of the faces of $\Gamma_{n}$. In [17], Yan, Yeung, and Zhang proved a formula involving $\Gamma_{n}^{*}$ for the capacity of multi-source multi-sink network coding. Those constraints in the formula induced by network topology and source independency form a face of $\Gamma_{n}$, which shows that this holy-grail network coding problem corresponds to the entropy function characterizations on this face. For the secret-sharing problem, see [18] for example, the perfect secrecy and decoding correctness conditions of an access structure determine a face $F$ on $\Gamma_{n}$, and the information ratio of the secret-sharing problem can be considered as an optimization problem whose feasible region is $F\cap\Gamma^{*}_{n}$. Other problems, such as distributed data storage [19], coded caching [20], Markov random fields [21], and relational database [22] are also related to the entropy regions on the faces of $\Gamma_{n}$. Recently, there are a series of works on how to compute the faces of $\Gamma_{n}$ more efficiently by machine [23, 24, 25, 26].
Though the information theory problems above usually involve more than four random variables, and the corresponding faces are of dimension higher than $2$, following the characterizations of extreme rays. i.e., $1$-dimensional faces containing matroid [7, 8, 9], and $2$-dimensional faces of $\Gamma_{3}$ [10, 11], in this series of two papers, we characterize entropy functions on the $2$-dimensional faces of $\Gamma_{4}$, which may serve as stepping stones to the general cases of this problem.
In Part I, we first characterize entropy functions on extreme rays containing integer polymatroids $\hat{U}_{2,5}$ and $\hat{U}_{3,5}$ (see Subsection II-B for the definition), which cannot be find in any existing literatures. We then enumerate all $59$ types of $2$-dimensional faces of $\Gamma_{4}$ by an algorithm and completely characterize $49$ types of them. Among them, $13$ types are all-entropic faces, while $7$ are non-entropic ones [^1] By [14, 15, 16], all $2$-dimensional faces containing Vámos ray $V_{8}$ are non-entropic, and they are also all of the $2$-dimensional faces crosscut by Zhang-Yeung, or any other non-Shannon-type inequalities.. The characterization of $17$ types can be obtained by the $2$ types of non-trivial faces of $\Gamma_{3}$ characterized in [10] and [11]. Of the $22$ types of non-trivial faces, $12$ will be characterized in Part I. The primary technique we adopt to characterize these non-trivial faces is the graph-coloring in [10]. For a $3$-dimensional random vector $(X_{i},i\in N_{3})$, if $X_{1}$ and $X_{2}$ are independent, and $X_{3}$ is a function of $(X_{1},X_{2})$, $X_{1}$ and $X_{2}$ can be considered distributed on the two sets of vertices $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ of a bipartite graph with edges those $(x_{1},x_{2})$ such that $p(x_{1},x_{2})>0$, and these edges are colored by $x_{3}\in\mathcal{X}_{3}$. In our work, when the number of the random variables becomes four, we color bipartite graphs by two color systems or color a tripartite graph by one color system.
The remaining $10$ cases of faces will be characterized in Part II [27] of the two-paper series. Among them, $8$ faces are fully and $2$ are partially characterized. This part has a more combinatorial flavour. All integer polymatroid with rank exceeding $1$ contained by an extreme ray will be associated a combinatorial structure. Then the entropy functions on $2$-dimensional faces containing $2$ such extreme rays breed new combinatorial-design structures, generalizing corresponding combinatorial structures in both extreme rays. These structures can also be considered as generalizations of $\rm VOA$s introduced in [7, 8].
The rest of this paper, i.e., Part I, is organized as follows. Section II gives the preliminaries on matroids, polyhedral cones, and the extreme rays of $\Gamma_{4}$. All the $59$ types of faces are generated by Algorithm I and listed in Table 3 in Section III. In Section IV, we fully characterize $49$ types among all the $2$-dimensional faces of $\Gamma_{4}$, and the results are listed in Table LABEL:long.
II Preliminaries
II-A Matroids
For a polymatroid $\mathbf{h}\in\Gamma_{n}$, if $\mathbf{h}(A)\in\mathbb{Z}$ for all $A\subseteq N_{n}$, $\mathbf{h}$ is called integer. An ordered pair $M=(N_{n},\mathbf{r})$ is called a matroid with rank function $\mathbf{r}$ if $\mathbf{r}$ is an integer polymatroid with $\mathbf{r}(A)\leq|A|$ for all $A\subseteq N_{n}$. Like polymatroids, in this paper, we do not distinguish a matroid and its rank function unless otherwise specified.
A uniform matroid $U_{k,n}$ is a matroid with rank function $\mathbf{r}(A)=\min\{k,|A|\}$ for all $A\subseteq N_{n}$.
For a matroid $M$ and $e\in N_{n}$, if $\mathbf{r}(e)=0$, $e$ is called a loop of $M$. For $e,e^{\prime}\in N_{n}$, if $\mathbf{r}(\{e,e^{\prime}\})=1$, then $e$ and $e^{\prime}$ are called parallel.
For more about matroid theory, readers can refer to [28].
II-B Faces of a polyhedral cone
Let $C\subseteq\mathbb{R}^{d}$ be a full-dimensional polyhedral cone. For a hyperplane $P$ containing the origin $O$ in $\mathbb{R}^{d}$, if $C\subseteq P^{+}$, where $P^{+}$ is one of the two halfspaces corresponding to $P$, $F\triangleq C\cap P$ is called a (proper) face of $C$, while $C$ itself is its improper face. When $\dim F=d-1$, $F$ is called a facet of $C$, and when $\dim F=1$, $F$ is an extreme ray of $C$. Either the set of all facets or the set of all extreme rays of $C$ uniquely determines the cone, and they are called H-representation and V-representation of the cone, respectively. Each face $F$ of the cone can be written as the intersection of the facets of the cone that contains $F$, or the convex hull of the extreme rays contained in $F$. We also call them the H-representation and V-representation of $F$, respectively. More details on this topic are referred to [29].
As discussed in Section I, $\Gamma_{n}$ is a polyhedral cone in $\mathcal{H}_{n}$ determined by polymatroidal axioms in (I.1)-(I.3). They are equivalent to the following elemental inequalities
$\mathbf{h}(N_{n})\geq\mathbf{h}({N_{n}\setminus\{i\}})\quad\quad i\in N_{n};$ (II.B.1)
$\mathbf{h}(K)+\mathbf{h}(K\cup ij)\leq\mathbf{h}(K\cup i)+\mathbf{h}(K\cup j),$ (II.B.2)
$\quad\quad i,j\in N_{n},K\subseteq N_{n}\setminus\{i,j\}$
each of which determines a facet of $\Gamma_{n}$ [2, Chapter 14]. When $n=4$, it can be checked that there are $28$ elemental inequalities, or $28$ facets of $\Gamma_{4}$.
It can be seen in [30] that there are $41$ extreme rays of $\Gamma_{4}$. Note that each extreme ray $E_{\mathbf{r}}$ of $\Gamma_{4}$ can be written in the form
$E_{\mathbf{r}}=\{a\mathbf{r}:a\geq 0\}$ (II.B.3)
where $\mathbf{r}$ is the minimal integer polymatroid in the ray, that is, an integer polymatroid such that $\mathbf{r}/t$ is not integer for any integer $t>1$. .
Therefore, in this paper, we refer to the minimal integer polymatroid $\mathbf{r}$ as the extreme ray $E_{\mathbf{r}}$ containing it unless otherwise specified. Note that some of these integer polymatroids are matroids. The $41$ extreme rays can be classified into the following $11$ types.
  • $U^{i}_{1,1}$, $i\in N_{4}$;
  • $U^{\alpha}_{1,2}$, $\alpha\subseteq N_{4}$, $|\alpha|=2$;
  • $U^{\alpha}_{1,3}$, $\alpha\subseteq N_{4}$, $|\alpha|=3$;
  • $U^{\alpha}_{2,3}$, $\alpha\subseteq N_{4}$, $|\alpha|=3$;
for $U^{\alpha}_{k,m}$ with $\alpha\subseteq N_{4}$ and $|\alpha|=m$, we mean a matroid on $N_{4}$ whose restriction on $\alpha$ is a $U_{k,m}$ and $e\in N_{4}\setminus\alpha$ being loops;
  • $U_{1,4}$;
  • $\mathcal{W}^{\alpha}_{2}$, $\alpha\subseteq N_{4}$, $|\alpha|=2$;
for $\mathcal{W}^{\alpha}_{2}$ with $\alpha\subseteq N_{4}$, $|\alpha|=2$, it is called a wheel matroid with order 22, [^2] We adopt the notation and terminology in [28, Section 8.4] and it is a matroid with two parallel elements in $\alpha$, and each element in $\alpha$ and the other two elements in $N_{4}$ form a $U_{2,3}$;
  • $U_{2,4}$;
  • $U_{3,4}$;
  • $\hat{U}^{i}_{2,5}$, $i\in N_{4}$;
for $\hat{U}^{i}_{2,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{2,5}$, [^3] See [31, Theorem 1.3.6] or [32, Theorem 4] for the definition of free expansion. and its rank function is defined by
$\mathbf{r}(A)=\begin{cases}\min\{2,|A|\}\quad&A\neq\{i\}\\ 2\quad&A=\{i\}\end{cases}$
for any $A\subseteq N_{4}$,
  • $\hat{U}^{i}_{3,5}$;
for $\hat{U}^{i}_{3,5}$ with $i\in N_{4}$, it is a polymatroid whose free expansion is a $U_{3,5}$, and its rank function is defined by
$\mathbf{r}(A)=\begin{cases}\min\{3,|A|+1\},\quad&i\in A,\\ |A|,\quad&i\not\in A,\end{cases}$
for any $A\subseteq N_{4}$;
  • $V^{\alpha}_{8}$, $\alpha\subseteq N_{4}$, $|\alpha|=2$;
for $V^{\alpha}_{8}$ with $\alpha\subseteq N_{4}$ and $|\alpha|=2$, it is a polymatroid whose free expansion is the Vámos matroid, and its rank function is defined by
$\mathbf{r}(A)=\begin{cases}3,\quad&|A|=2\text{ and }A\neq\alpha,\\ \min\{4,2|A|\},\quad&\text{ o.w.}\end{cases}$
It can be seen that for an extreme ray in the form $E^{\alpha}$ with $\alpha\subseteq N_{4}$, it is in a type with $\binom{4}{|\alpha|}$ extreme rays, and each extreme ray in the type can be obtained from each other by permuting the indices in $N_{4}$. For a specific extreme ray, say $U^{\{1,2\}}_{1,2}$, for simplicity, we will drop the bracket and comma of the set in the superscript, and write it as $U^{12}_{1,2}$. To facilitate the readers, the $11$ types of extreme rays are presented in Table 1 by the rank functions of their minimal integer polymatroid in the form of $15$-dimensional vectors, i.e., $\mathbf{r}=(\mathbf{r}(A),\emptyset\neq A\subseteq N_{4})$.
In the same manner, we denote the $8$ extreme rays of $\Gamma_{3}$ and classify them in the following $4$ types.
  • $U^{i,3}_{1,1}$, $i\in N_{3}$;
  • $U^{\alpha,3}_{1,2}$, $\alpha\subseteq N_{3}$, $|\alpha|=2$;
  • $U_{1,3}$;
  • $U_{2,3}$.
Here we put second superscript $3$ to $U^{i,3}_{1,1}$ and $U^{\alpha,3}_{1,1}$ to distinguish them from $U^{i}_{1,1}$ and $U^{\alpha}_{1,1}$, the extreme rays of $\Gamma_{4}$, respectively.
Table 1: Extreme rays of $\Gamma_{4}$ and their rank functions
[…]
$U^{1}_{1,1}$ 11 0 0 0 1 11 11 0 0 0 11 11 11 11 11
$U^{12}_{1,2}$ 11 11 0 0 11 11 11 11 11 0 11 11 11 11 11
$U^{123}_{1,3}$ 11 11 11 0 11 11 11 11 11 11 11 11 11 11 11
$U_{1,4}$ 11 11 11 11 11 11 11 11 11 11 11 11 11 11 11
$U^{123}_{2,3}$ 11 11 11 0 22 22 11 22 11 11 22 22 22 22 22
$\mathcal{W}^{12}_{2}$ 11 11 11 11 11 22 22 22 22 22 22 22 22 22 22
$U_{2,4}$ 11 11 11 11 22 22 22 22 22 22 22 22 22 22 22
$U_{3,4}$ 11 11 11 11 22 22 22 22 22 22 33 33 33 33 33
$\hat{U}^{1}_{2,5}$ 22 11 11 11 22 22 22 22 22 22 22 22 22 22 22
$\hat{U}^{1}_{3,5}$ 22 11 11 11 33 33 33 22 22 22 33 33 33 33 33
$V^{12}_{8}$ 22 22 22 22 44 33 33 33 33 33 44 44 44 44 44
II-C Entropy functions on the extreme rays of $\Gamma_{4}$
As discussed in Subsection II-B, the V-representation of a face can be written as the convex hull of the extreme rays it contains. Therefore, to characterize entropy functions on $2$-dimensional faces of $\Gamma_{4}$, we need first to characterize the entropy functions on the extreme rays of $\Gamma_{4}$. The results of the characterizations of all extreme rays in this subsection are summarized in Table 2.
In the following, extreme rays are written in the form in (II.B.3). For each type of extreme ray, we only consider one of them. Throughout this paper, for any random vector $(X_{i}:i\in N_{4})$, each $X_{i}$ is assumed distributed on a finite set $\mathcal{X}_{i},i\in N_{4}$. For any marginal distribution of $(X_{i}:i\in N_{4})$, say $p_{X_{2}X_{3}}(x_{2},x_{3})$ with $(x_{2},x_{3})\in\mathcal{X}_{2}\times\mathcal{X}_{3}$, we usually drop its subscription and write it as $p(x_{2},x_{3})$ if there is no ambiguity. We may further simplify it as $p(x_{23})$.
Theorem II-C.1.
For $E=U^{1}_{1,1},U^{12}_{1,2},U^{123}_{1,3},U_{1,4}$, i.e., extreme rays containing a matroid with rank $1$, $a\mathbf{r}\in E$ is entropic for all $a\geq 0$.
Proof.
Let $X_{1}$ be any random variable with $H(X_{1})=a$. Let $X_{i}=X_{1}$ for any $i\in\alpha$ and $X_{j}$ is a constant if $j\in N_{4}\setminus\alpha$, where $\alpha=N_{k}$ if $E=U^{\alpha}_{1,k},k=1,2,3,4$ and here $U^{N_{k}}_{1,4}=U_{1,4}$. It is readily seen that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $a\mathbf{h}$. ∎
The characterization of the following four types of extreme rays follows immediately from matroidal entropy functions in [7] and
Theorem II-C.2.
For $E=U^{123}_{2,3},\mathcal{W}^{14}_{2},U_{3,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log v$ for integer $v\geq 1$.
Theorem II-C.3.
([7, Proposition 2]) For $E=U_{2,4}$, $a\mathbf{r}\in E$ is entropic if and only if $a=\log v$ for positive integer $v\neq 2,6$.
Table 2: Entropy functions on the extreme rays of $\Gamma_{4}$
Extreme rays $E$ Theorems Figures of entropy region $E^{*}=E\cap\Gamma^{*}_{4}$ Entropy region $E^{*}$
$U^{i}_{1,1}$ II-C.1 [Uncaptioned image] $\{a\mathbf{r}:a\geq 0\}$
$U^{\alpha}_{1,2}$ II-C.1 [Uncaptioned image] $\{a\mathbf{r}:a\geq 0\}$
$U^{\alpha}_{1,3}$ II-C.1 [Uncaptioned image] $\{a\mathbf{r}:a\geq 0\}$
$U_{1,4}$ II-C.1 [Uncaptioned image] $\{a\mathbf{r}:a\geq 0\}$
$U^{\alpha}_{2,3}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\mathcal{W}^{\alpha}_{2}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$U_{2,4}$ II-C.3 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,v>0,v\in\mathbb{Z},v\neq 2,6\}$
$U_{3,4}$ II-C.2 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\hat{U}^{i}_{2,5}$ II-C.4 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$\hat{U}^{i}_{3,5}$ II-C.5 [Uncaptioned image] $\{a\mathbf{r}:a=\log v,\ v>0,v\in\mathbb{Z}\}$
$V^{\alpha}_{8}$ II-C.6 [Uncaptioned image] $\{a\mathbf{r}:a=0\}$
It is well known that the rays containing a minimal integer polymatroid with rank-$1$ are all-entropic. It has been proven that the vectors $a\mathbf{r}$ in $U^{\alpha}_{2,3}$ are entropic if and only if $a=\log{v}$ for $k\in\mathbb{Z}$ [2, Theorem 15.2]. In the ray $U_{2,4}$, $a$ can take the values $\log{v}$ except for $\log 2$ and $\log 6$ [8, Proposition 6]. In the ray $U_{3,4}$, the region is the same as in $U^{\alpha}_{2,3}$. These three cases correspond to the Latin Square, Orthogonal Two Latin Squares, and Latin Cube, respectively. As for the ray $\hat{U}^{i}_{2,5}$ and $\hat{U}^{i}_{3,5}$, they are implied by the following theorems. For the ray $V^{\alpha}_{8}$, the fact that it disobeys the Zhang-Yeung inequality indicates that it lies outside of $\overline{\Gamma_{4}^{*}}$ and, consequently, is non-entropic.
Theorem II-C.4.
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{2,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log v$ for integer $v>0$.
Proof.
For $\mathbf{h}=a\mathbf{r}$, where $\mathbf{r}$ is the rank function of $\hat{U}^{1}_{2,5}$, if we restrict it on $\{2,3,4\}$, we will see that its restriction $\mathbf{h}^{\prime}=a\mathbf{r}_{1}$, where $\mathbf{r}_{1}$ is the rank function of $U_{2,3}$, is entropic only if $a=\log v$ for integer $v>0$.
To prove the “if” part, let $X_{2}$ and $X_{3}$ be independent and uniformly distributed on $\mathbb{I}_{v}\triangleq\{0,\dots,v-1\}$, and $X_{4}=X_{2}+X_{3}\mod v$. Let $X_{1}=v\cdot X_{2}+X_{3}$. Then $X_{1}$ is uniformly distributed on $\mathbb{I}_{v^{2}}$ and it can be checked that $(X_{i},i\in N_{4})$ such constructed characterizes $\mathbf{h}=\log{v}\cdot\mathbf{r}$. ∎
Lemma 1.
Assume that each $X_{i},i\in N_{3}$ is distributed on $\mathcal{X}_{i}$ and for $x_{i}\in\mathcal{X}_{i},p(x_{i})>0$. If $X_{1}$ and $X_{2}$ are independent and for any $p(x_{1},x_{2},x_{3})>0$, $p(x_{1})=p(x_{2})$, then $X_{1}$ and $X_{2}$ are uniformly distributed on $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$, respectively, $|\mathcal{X}_{1}|=|\mathcal{X}_{2}|$ and $H(X_{1})=H(X_{2})$.
Proof.
Let $X_{1}$ and $X_{2}$ be independent and $p(x_{1},x_{2},x_{3})>0$. By assumption, we have
$p(x_{1})=p(x_{2}).$ (II.C.1)
Since $X_{1}$ and $X_{2}$ are independent, for any $x_{1}^{\prime}\in\mathcal{X}_{1}$ with $x_{1}^{\prime}\neq x_{1}$, we obtain
$p(x_{1}^{\prime},x_{2})=p(x_{1}^{\prime})p(x_{2})>0.$
As
$p(x_{1}^{\prime},x_{2})=\sum_{x_{3}}p(x_{1}^{\prime},x_{2},x_{3}),$
there exists $x_{3}^{\prime}\in\mathcal{X}_{3}$ such that $p(x_{1}^{\prime},x_{2},x_{3}^{\prime})>0$, which implies
$p(x_{1}^{\prime})=p(x_{2})$ (II.C.2)
according to the assumption. By (II.C.1) and (II.C.2),
$p(x_{1})=p(x_{1}^{\prime}),$ (II.C.3)
which implies $p(x_{1})$ are equal for all $x_{1}\in\mathcal{X}_{1}$. Then, we conclude that $X_{1}$ is uniformly distributed on $\mathcal{X}_{1}$. By symmetry, $X_{2}$ is uniformly distributed on $\mathcal{X}_{2}$. As $p(x_{1})=p(x_{2})$ for all $x_{1}\in\mathcal{X}_{1}$ and $x_{2}\in\mathcal{X}_{2}$, then $|\mathcal{X}_{1}|=|\mathcal{X}_{2}|$ and $H(X_{1})=H(X_{2})$. ∎
Theorem II-C.5.
For the rank function $\mathbf{r}$ of $\hat{U}^{i}_{3,5}$, $\mathbf{h}=a\mathbf{r}$ is entropic if and only if $a=\log v$ for integer $v>0$.
Proof.
If $\mathbf{h}\in F=\hat{U}^{1}_{3,5}$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i\neq j,i,j\in N_{4}$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{i,j,k}),$
$\qquad\qquad\qquad\quad\ \{i,j,k\}=\{2,3,4\}$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\qquad\qquad|K|=2,K\nsubseteq\{2,3,4\}.$
Assume that each $X_{i}$ is distributed on $\mathcal{X}_{i}$ and for $x_{i}\in\mathcal{X}_{i}$, $p(x_{i})>0$. For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (II.C.4)
$=p(x_{1},x_{2},x_{4})$ (II.C.5)
$=p(x_{1},x_{3},x_{4})$ (II.C.6)
$=p(x_{2},x_{3},x_{4}),$ (II.C.7)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (II.C.8)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (II.C.9)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (II.C.10)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (II.C.11)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (II.C.12)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (II.C.13)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (II.C.14)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (II.C.15)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (II.C.16)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (II.C.17)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (II.C.18)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (II.C.19)
By (II.C.4)-(II.C.7) and (II.C.17)-(II.C.19), we have
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (II.C.20)
$=p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4}).$ (II.C.21)
In light of (II.C.8)-(II.C.10), relapcing $p(x_{1},x_{2}),p(x_{1},x_{3})$ and $p(x_{1},x_{4})$ by $p(x_{1})p(x_{2}),p(x_{1})p(x_{3})$ and $p(x_{1})p(x_{4})$ in (II.C.21), we obtain
$p(x_{2})=p(x_{3})=p(x_{4}).$ (II.C.22)
By Lemma 1, $X_{2},X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{2},\mathcal{X}_{3}$ and $\mathcal{X}_{4}$ and $H(X_{2})=H(X_{3})=H(X_{4})=\log v$, where $v=|\mathcal{X}_{i}|,i\in\{2,3,4\}$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have $H(X_{2})=H(X_{3})=H(X_{4})=a$. Thus $a$ can only take the value $\log v$.
Now we prove that all of these polymatroids are entropic. For $a=\log v$, let $X_{2},X_{3}$ and $X_{4}$ be mutually independent and uniformly distributed on $\mathbb{I}_{v}$. Let
$X_{1}=v$ $[(X_{3}-X_{2})\mod v]$
$+$ $[(X_{4}-X_{2})\mod v].$ (II.C.23)
It remains to check the entropy function of such constructed $(X_{i},i\in N_{4})$ is $a\mathbf{r}$. Given that $X_{2}$, $X_{3}$ and $X_{4}$ are mutually independent and uniformly distributed, we have
$H(X_{2})$ $=H(X_{3})=H(X_{4})=\log v$ (II.C.24)
$H(X_{234})$ $=H(X_{2})+H(X_{3})+H(X_{4})$ (II.C.25)
$=3\log v.$ (II.C.26)
Then by (II.C.23), we obtain $H(X_{1234})=H(X_{234})$. Dividing both sides of (II.C.23) by $v$ and then rounding them down, we obtain
$\lfloor\dfrac{X_{1}}{v}\rfloor=(X_{3}-X_{2})\mod v$ (II.C.27)
Thus
$X_{3}\mod v=(X_{2}+\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v.$ (II.C.28)
Note that $X_{3}$ is distributed on $\mathbb{I}_{v}$, we have
$X_{3}=(X_{2}+\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v,$ (II.C.29)
which implies that $X_{3}$ is a function of $(X_{1},X_{2})$. By the same argument,
$X_{2}=(X_{3}-\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v.$ (II.C.30)
By (II.C.27), replacing $(X_{3}-X_{2})\mod v$ by $\lfloor\dfrac{X_{1}}{v}\rfloor$ in (II.C.23),
$X_{1}=v\lfloor\dfrac{X_{1}}{v}\rfloor+[(X_{4}-X_{2})\mod v].$ (II.C.31)
Similarly,
$X_{4}$ $=(X_{1}+X_{2}-\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v,$ (II.C.32)
$X_{2}$ $=(X_{4}-X_{1}+\lfloor\dfrac{X_{1}}{v}\rfloor)\mod v$ (II.C.33)
Based on (II.C.29), (II.C.30), (II.C.32) and (II.C.33), the random vetor $(X_{i},i\in N_{4})$ must satisfy the following equalities,
$H(X_{123})$ $=H(X_{12})=H(X_{13}),$ (II.C.34)
$H(X_{124})$ $=H(X_{12})=H(X_{14}).$ (II.C.35)
Note that
$I(X_{3};X_{4}|X_{12})=$ $H(X_{123})+H(X_{124})$
$-H(X_{12})-H(X_{1234}).$ (II.C.36)
In light of (II.C.34) and (II.C.35), replacing $H(X_{123})$ and $H(X_{124})$ by $H(X_{12})$ in (II.C.36),
$I(X_{3};X_{4}|X_{12})$ $=H(X_{12})-H(X_{1234})$ (II.C.37)
$=-H(X_{34}|X_{12})$ (II.C.38)
By the nonnegativity of $H(X_{34}|X_{12})$ and $I(X_{3};X_{4}|X_{12})$, we have
$H(X_{1234})=H(X_{12}).$ (II.C.39)
Together with (II.C.23), which implies
$H(X_{1234})=H(X_{234}),$ (II.C.40)
and (II.C.26), we have
$H(X_{12})=3\log v.$ (II.C.41)
By (II.C.39) and (II.C.34), $3\log v=$ $H(X_{13})$$\leq H(X_{134})\leq H(X_{1234})=3\log v$ which implies
$H(X_{134})=3\log v.$ (II.C.42)
The mutual independence of $X_{2}$, $X_{3}$ and $X_{4}$ and (II.C.24) implies that
$H(X_{23})=H(X_{24})=H(X_{34})=2\log v$ (II.C.43)
By (II.C.24), (II.C.41), (II.C.34), (II.C.35), (II.C.43), (II.C.39), (II.C.40) and (II.C.42) the value of $H(X_{A})$ are determined for all $A\subseteq N$ except for $A=\{1\}$. Now we verify that $H(X_{1})=2\log v$. By (II.C.24) and (II.C.41),
$H(X_{1})\geq H(X_{12})-H(X_{2})=2\log v.$ (II.C.44)
On the other hand, as $X_{1}$ is distributed on $\mathbb{I}_{v^{2}}$ by (II.C.23), we have $H(X_{1})\leq 2\log v$. The proof is accomplished. ∎
Theorem II-C.6.
For $E=V^{\alpha}$, $a\mathbf{h}\in E$ is entropic if and only if $a=0$.
Proof.
For $a=0$, $a\mathbf{h}$ is the origin, and it is characterized by constant $(X_{i},i\in N_{4})$. For $a>0$, all $a\mathbf{h}$ dissatisfy Zhang-Yeung inequality, and are so non-entropic. ∎
III Two-dimensional faces of $\Gamma_{4}$
As we discussed in the previous section, each face of a polyhedral cone has two representations, i.e, H-representation, the intersection of the facets containing the face, and V-representation, the convex hull of the extreme rays contained in the face. For $2$-dimensional faces of $\Gamma_{4}$, each of them contains exactly two extreme rays. Thus in this paper we adopt the V-representation. For each two extreme rays of $\Gamma_{4}$, to determine whether the convex hull of them forms a face of $\Gamma_{4}$, we only need to check that the intersection of the facets containing the two extreme rays is a face of dimension $2$, that is, containing no other extreme rays. The catalogue of all $2$-dimensional faces of $\Gamma_{4}$ is generated by Algorithm 1.
Algorithm 1 Two-dimensional Face of $\Gamma_{4}$ Generating Algorithm.
1:The family $\mathcal{F}$ of all $28$ facets and the family $\mathcal{E}$ of all 41 extreme rays of $\Gamma_{4}$.
2:Upper triangle of a $41\times 41$ $(0,1)$-matrix $C$,where $C(i,j)=1$ if and only if the convex hull of the $i$-th extreme ray $E_{i}$ and the $j$-th extreme ray $E_{j}$ forms a $2$-dimensional face of $\Gamma_{4}$.
3:for $1\leq i<j\leq 41$ do
4:  $C(i,j)\leftarrow 1$
5:  for $k=1$ to $28$ do
6:   if the $k$-th facet $F_{k}$ contains both $E_{i}$ and $E_{j}$, then
7:     put $F_{k}$ in $\mathcal{F}^{\prime}$.
8:   end if
9:  end for
10:  for $E\in\mathcal{E}\setminus\{E_{i},E_{j}\}$ do
11:   if $E$ is contained in the face $\cap_{F\in\mathcal{F}^{\prime}}F$ then
12:     $C(i,j)\leftarrow 0$
13:     break
14:   end if
15:  end for
16:end for
The result of Algorithm 1, i.e., $510$ 22-dimensional faces corresponding to those “11”-entries in $C$ are classified in Table 3. Similar to the fact that the $41$ extreme rays can be classified in $11$ types, the $510$ 22-dimensional faces can also be classified to $59$ types, and in each type, each face can be obtained from another by permuting the indices in $N_{4}$. These $59$ types of faces are listed in the upper triangle of Table 3. We label the rows and columns with the $11$ types of extreme rays of $\Gamma_{4}$. For simplicity, we denote the face $F=\mathrm{cone}(E_{1},E_{2})$ by $(E_{1},E_{2})$. In each cell with $(E_{1},E_{2})\ n$, Thm. mm”, $(E_{1},E_{2})$” denote a representative of the type of $2$-dimensional faces, where $E_{1}$ $(E_{2})$” is a representative of the type the extrem rays in the column (row), “nn” is the number of the faces in this type and this face is characterized in “Thm. mm” in Section IV. For the cell with “0”, the convex hull of the two extreme rays in each type forms no $2$-dimensional faces of $\Gamma_{4}$. More details of the faces in Table 3 are discussed in Section IV.
Table 3: Two-dimensional faces of $\Gamma_{4}$
$U^{i}_{1,1}$ $U^{\alpha}_{1,2}$ $U^{\alpha}_{1,3}$ $U_{1,4}$ $U^{\alpha}_{2,3}$ $\mathcal{W}^{\alpha}_{2}$ $U_{2,4}$ $U_{3,4}$ $\hat{U}^{i}_{2,5}$ $\hat{U}^{i}_{3,5}$ $V^{\alpha}_{8}$
$U^{j}_{1,1}$ $({U^{12}_{1,2},U^{1}_{1,1})}$ $({U^{123}_{1,3},U^{1}_{1,1})}$ $(U_{2,3}^{123},U_{1,1}^{1})$ $(\mathcal{W}_{2}^{14},U_{1,1}^{1})$ $(\hat{U}^{1}_{2,5},U_{1,1}^{1})$ $(\hat{U}^{1}_{3,5},U^{1}_{1,1})$ $(V^{12}_{8},U^{1}_{1,1})$
$({U^{1}_{1,1},U^{2}_{1,1})}$ $12$,Thm.IV-A.1 $12$, Thm.IV-A.1 $({U_{1,4},U^{1}_{1,1})}$ $12$, Thm.IV-B.2 $12$, Thm.IV-B.2 $(U_{2,4},U^{1}_{1,1})$ $(U_{3,4},U^{1}_{1,1})$ 44,Thm.IV-B.3 44,Thm.IV-B.5 $12$,Thm.IV-G.1
66, Thm.IV-A.1 $({U^{12}_{1,2},U^{3}_{1,1})}$ $({U^{123}_{1,3},U^{4}_{1,1})}$ 44, Thm.IV-A.1 $(U_{2,3}^{123},U_{1,1}^{4})$ $(\mathcal{W}_{2}^{34},U_{1,1}^{1})$ 44 44, Thm.IV-E.1 $(\hat{U}^{1}_{2,5},U_{1,1}^{2})$ $(\hat{U}^{1}_{3,5},U^{2}_{1,1})$ $(V^{12}_{8},U^{3}_{1,1})$
$12$, Thm.IV-A.1 44, Thm.IV-A.1 44, Thm.IV-B.2 $12$, Thm.IV-B.2 $12$, Thm.IV-B.3 $12$, Thm.IV-B.5 $12$, Thm.IV-G.1
$U^{\beta}_{1,2}$ $\backslash$ $(\mathcal{W}_{2}^{14},U_{1,2}^{14})$
$(U^{12}_{1,2},U^{13}_{1,2})$ $({U^{123}_{1,3},U^{12}_{1,2})}$ $(U_{2,3}^{123},U_{1,2}^{12})$ 66, Thm.IV-B.2
$12$, Thm.IV-A.1 $12$, Thm.IV-A.1 $({U_{1,4},U^{12}_{1,2})}$ $12$, Thm.IV-B.1 $(\mathcal{W}_{2}^{24},U_{1,2}^{14})$ $(U_{2,4},U^{12}_{1,2})$ $(U_{3,4},U^{12}_{1,2})$ $(\hat{U}^{1}_{2,5},U_{1,2}^{12})$ $(\hat{U}^{1}_{3,5},U^{12}_{1,2})$ $(V^{12}_{8},U^{13}_{1,2})$
$(U^{12}_{1,2},U^{34}_{1,2})$ $({U^{123}_{1,3},U^{14}_{1,2})}$ 66, Thm.IV-A.1 $(U_{2,3}^{123},U_{1,2}^{14})$ $24$, Thm.IV-B.2 66 66, Thm.IV-E.2 $12$, Thm.IV-B.3 $12$, Thm.IV-B.5 $24$, Thm.IV-G.1
33, Thm.IV-A.1 $12$, Thm.IV-A.1 $12$, Thm.IV-B.2 $(\mathcal{W}_{2}^{34},U_{1,2}^{12})$
66, Thm.IV-B.1
[…]
66, Thm.IV-A.1 44, Thm.IV-A.1 $12$, Thm.IV-C.1 $12$, Thm.IV-B.1 44 44, Thm.IV-E.4 $12$, Thm.IV-B.4 44, Thm.IV-F.1 $12$, Thm.IV-G.1
$U_{1,4}$ $\backslash \\backslash \\$backslash $(U_{2,3}^{123},U_{1,4})$ 0 0 $(U_{3,4},U_{1,4})$ 0 0 $(V^{12}_{8},U_{1,4})$
44, Thm.IV-C.3 11, Thm.IV-E.5 66, Thm.IV-G.1
$U^{\beta}_{2,3}$ $\backslash \\backslash \\$backslash $(U_{2,3}^{123},U_{2,3}^{124})$ $(\mathcal{W}_{2}^{12},U_{2,3}^{134})$ $(U_{2,4},U^{123}_{2,3}),4$ $(U_{3,4},U^{123}_{2,3})$ $(\hat{U}^{1}_{2,5},U_{2,3}^{234})$ $(\hat{U}^{1}_{3,5},U^{123}_{2,3})$ $(V^{12}_{8},U^{123}_{2,3})$
66, Thm.IV-C.2 $12$, Thm.IV-D.1 44 44, Thm.IV-E.3 44 $12$, Thm.IV-F.2 $12$, Thm.IV-G.1
$\mathcal{W}^{\beta}_{2}$ $\backslash \\backslash \\$backslash $\backslash$ $(\mathcal{W}_{2}^{12},\mathcal{W}_{2}^{13})$ $(U_{2,4},\mathcal{W}^{12}_{2})$ 0 $(\hat{U}^{1}_{2,5},\mathcal{W}_{2}^{12})$ $(\hat{U}^{1}_{3,5},\mathcal{W}^{23}_{2})$ 0
$12$ 66 $12$ $12$, Thm.IV-F.3
$U_{2,4}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash 0 $(\hat{U}^{1}_{2,5},U_{2,4})$ $(\hat{U}^{1}_{3,5},U_{2,4})$ 0
44 44
$U_{3,4}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash $\backslash$ 0 0 $(V^{12}_{8},U_{3,4})$
66, Thm.IV-G.1
$\hat{U}^{j}_{2,5}$ $\backslash \\backslash \\$$\backslash \\backslash \\$backslash $\backslash$ 0 0 0
$\hat{U}^{j}_{3,5}$ $\backslash \\backslash \\$$\backslash \\backslash \\$$\backslash \\backslash$ 0 0
$V^{\beta}_{8}$ $\backslash \\backslash \\$$\backslash \\backslash \\$$\backslash \\backslash \\$backslash 0
IV Characterization of entropy functions on two-dimensional faces of $\Gamma_{4}$
In this section, we characterize the entropy functions on $2$-dimensional faces of $\Gamma_{4}$. We embed each face $F=(E_{1},E_{2})$ in the first octant of a $2$-dimensional cartesian coordinate system whose axis are labeled by $a$ and $b$. Thus, for each $(a,b)$, $a,b\geq 0$, it represents the polymatroid $a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{i},i=1,2$ is the rank function of the minimal integer polymatroid in $E_{i}$, respectively. Throughout this paper, for a random vector $(X_{i},i\in N_{4})$ or its subvectors, we assume each $X_{i}$ is distributed on a finite set $\mathcal{X}_{i}$, and for each $x_{i}\in$ $\mathcal{X}_{i}$, $p(x_{i})>0$.
Lemma 2 (Lemma 15.3, [2]).
For any $\mathbf{h}_{1},\mathbf{h}_{2}\in\Gamma^{*}_{n}$, $\mathbf{h}_{1}+\mathbf{h}_{2}\in\Gamma^{*}_{n}$.
Lemma 3.
For a random vector $X_{i},i\in N_{3}$, consider the tripartite graph $G=(V,E)$, where $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}$ and $(x_{i},x_{j})\in E$ if $p(x_{i},x_{j})>0$ for distinct $i,j\in N_{3}$. If $(X_{i},i\in N_{3})$ satisfies the following information equalities,
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),$
where $i,j,k$ be any permutation of $1,2,3$, then each connected component of $G$ is a complete tripartite graph. Futhermore, if $p(x_{1})=p(x_{2})=p(x_{3})$ holds for any $p(x_{1},x_{2},x_{3})>0$, then the number of vertices in $\mathcal{X}_{i}$, $i=1,2,3$ are the same and the the probability mass of all of the vertices, the edges and the triangles are equal, respectively, in each connected component.
Proof.
For $(x_{i},i\in N_{3})\in\mathcal{X}_{N_{3}}$ with $p(x_{123})>0$, the information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (IV..1)
$p(x_{1},x_{2})p(x_{2},x_{3})$ $=p(x_{2})p(x_{1},x_{2},x_{3}),$ (IV..2)
$p(x_{1},x_{3})p(x_{2},x_{3})$ $=p(x_{3})p(x_{1},x_{2},x_{3}).$ (IV..3)
To prove the lemma, for such $(X_{i},i\in N_{3})$, we now prove the following two claims.
Claim 1.
For any two vertices $x_{1}\in\mathcal{X}_{1},x_{2}\in\mathcal{X}_{2}$, if they are adjacent to the same vertex in $\mathcal{X}_{3}$, then these two vertices are also adjacent.
Proof of Claim 1.
For any $x_{1}\in\mathcal{X}_{1}$, $x_{2}\in\mathcal{X}_{2}$ adjacent to the same vertex $x_{3}\in\mathcal{X}_{3}$, by definition of $G$, $p(x_{1},x_{3})$ and $p(x_{2},x_{3})$ are positive. By (IV..3), we have $p(x_{1},x_{2},x_{3})>0$ which implies $p(x_{1},x_{2})>0$ and so $x_{1}$ is adjacent to $x_{2}$. ∎
Claim 2.
For any two vertices $x_{1},x^{\prime}_{1}\in\mathcal{X}_{1}$, if they are adjacent to a common vertex in $\mathcal{X}_{2}$ or $\mathcal{X}_{3}$, then each $y\in\mathcal{X}_{2}$ or $\mathcal{X}_{3}$, either it is adjacent to both $x_{1}$ and $x^{\prime}_{1}$, or it is adjacent to none of them.
Proof of Claim 2.
Assume there exists two vertices $x_{1}$, $x_{1}^{\prime}\in\mathcal{X}_{1}$ with $x_{1}\neq x_{1}^{\prime}$ and they are adjacent to $x_{2}\in\mathcal{X}_{2}$, then $p(x_{1},x_{2})>0$ and $p(x_{1}^{\prime},x_{2})>0$. As
$p(x_{1},x_{2})=\sum_{x_{3}}p(x_{1},x_{2},x_{3})$
there exists $x_{3}\in\mathcal{X}_{3}$ such that $p(x_{1},x_{2},x_{3})>0$. Thus both $x_{1}$ and $x_{2}$ are adjacent to $x_{3}$. Since $x_{1}^{\prime}$ is adjacent to $x_{2}$, by Claim 1, $x_{1}^{\prime}$ is adjacent to $x_{3}$. For any vertex $y$ adjacent to $x_{1}^{\prime}$, by Claim 1, $y$ is adjacent to $x_{2}$ or $x_{3}$. As $x_{1}$ is adjacent to $x_{2}$ and $x_{3}$, we conclude that $y$ is adjacent to $x_{1}$. By symmetry, $x_{1}^{\prime}$ is adjacent to all vertices adjacent to $x_{1}$. ∎
Now we prove that each connected component is a complete tripartite graph. In a connected component, for any two vertices $x_{i}\in\mathcal{X}_{i},x_{j}\in\mathcal{X}_{j}$, $i,j\in N_{3},i\neq j$, there exists a path from $x_{i}$ to $x_{j}$ and the length of the path is denoted by $n$. We prove that $x_{i}$ and $x_{j}$ are adjacent by induction on $n$. First, it is obviously ture for $n=1$ and it holds for $n=2$ by Claim 1. Assume it is ture for some positive integers k1k-1 and $k$ where $k\geq 2$. Now consider $n=k+1$. If the vertex $y$ adjacent to $x_{j}$ on the path is not in $\mathcal{X}_{i}$, by induction hypothesis on $k$, we see that $x_{i}$ is adjacent to $y$. Then by Lemma 1, $x_{i}$ is adjacent to $x_{j}$. If $y$ is in $\mathcal{X}_{i}$, the vertex $z(z\neq x_{j})$ adjacent to $y$ on the path is adjacent to $x_{i}$ by induction on k1k-1. Then by Claim 2, $x_{i}$ is adjacent to $x_{j}$.
For $p(x_{1},x_{2},x_{3})>0$, if the probability mass function satisfies $p(x_{1})=p(x_{2})=p(x_{3})$, then by (IV..1)-(IV..3),
$p(x_{1},x_{2})=p(x_{1},x_{3})=(x_{2},x_{3}),$ (IV..4)
which implies that the probability mass of vertices and edges are equal in each triangle. Since each connected component is a complete tripartite graph, we conclude that the probability of vertex and edge are equal in each connected component. Then by (IV..1), the probability of the triangles are all equal in each connected component. The proof is accomplished. ∎
IV-A All-entropic faces
Theorem IV-A.1.
For $F=(E_{1},E_{2})$, where distinct $E_{i},i=1,2$ contains a rank-$1$ matroid, i.e., the first $4$ types of extreme rays, any $\mathbf{h}\in F$ is entropic.
Proof.
The fact that the whole $E_{i}$ is entropic, together with Lemma 2 immediately implies the theorem. ∎
Theorem IV-A.1 is also a corollary of [33, Proposition 4] or [34, Theorem 2].
IV-B Entropy functions on the faces extended from $2$-dimensional faces of $\Gamma_{3}$
In this subsection, we characterize those faces which can be obtained from the characterizations of $2$-dimensional faces of $\Gamma_{3}$. Thm.IV-B.1 characterize $3$ faces in Table 3 which can be obtained by the characterization of $2$-dimensional face $(U_{2,3},U^{12,3}_{1,2})$ of $\Gamma_{3}$. It has been done in [10] . Thus, the entropy functions on them have the same shape as those on $(U_{2,3},U^{12,3}_{1,2})$, which is depicted in Figure 1.
Theorem IV-B.1.
For $F=(U_{2,3}^{123},U_{1,2}^{12}),(\mathcal{W}_{2}^{34},U_{1,2}^{12})$, and $(\mathcal{W}_{2}^{14},U_{1,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b\geq\log\lceil 2^{a}\rceil$. [^4] We take the base of the logarithm be $2$ for computing the entropy.
Proof.
Let $F$ be any of the $3$ faces and $\mathbf{h}\in F$. Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the matroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i},i=1,2$ are the restriction of $\mathbf{r}_{i}$ on $N_{3}$. It can be checked that they are the rank functions of $U_{2,3}$ and $U^{12,3}_{1,2}$, respectively. Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{12,3}_{1,2})$ and $a+b\geq\log\lceil 2^{a}\rceil$ by [10, Theorem 1].
Now we prove the “if ” part of the theorem. Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}\in F$ and $a+b\geq\log\lceil 2^{a}\rceil$. Let $\mathbf{h}^{\prime}$ be the restriction of $\mathbf{h}$ on $N_{3}$. Therefore, $\mathbf{h}^{\prime}\in(U_{2,3},U^{12,3}_{1,2})$ and let $(X_{1},X_{2},X_{3})$ be its characterizing random vector. Now for
  • $F=(U_{2,3}^{123},U_{1,2}^{12})$, let $X_{4}$ be a constant;
  • $F=(\mathcal{W}_{2}^{34},U_{1,2}^{12})$, let $X_{4}=X_{3}$;
  • $F=(\mathcal{W}_{2}^{14},U_{1,3}^{124})$, let $X_{4}=X_{1}$.
Then $(X_{i},i\in N_{4})$ is the characterizing random vector of $\mathbf{h}$, which implies that $\mathbf{h}$ is entropic. ∎
Figure 1: The faces $(U_{2,3}^{123},U_{1,2}^{12})$, $(\mathcal{W}_{2}^{34},U_{1,2}^{12})$ and $(\mathcal{W}_{2}^{14},U_{1,3}^{124})$
For the $7$ faces characterized in Thm.IV-B.2, the entropy region on them have the same shape as $2$-dimensional face $(U_{2,3},U^{1,3}_{1,1})$ of $\Gamma_{3}$, which has been characterized in [11].
Theorem IV-B.2.
For $F=(U_{2,3}^{123},U_{1,1}^{1})$, $(U_{2,3}^{123},U_{1,1}^{4})$, $(U_{2,3}^{123},U_{1,2}^{14})$, $(\mathcal{W}_{2}^{14},U_{1,1}^{1})$, $(\mathcal{W}_{2}^{34},U_{1,1}^{1})$, $(\mathcal{W}_{2}^{14},U_{1,2}^{14})$ and $(\mathcal{W}_{2}^{24},U_{1,2}^{14})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{v}$ for integer $v>0$.
[…]
Proof.
For $F=(U_{2,3}^{123},U_{1,1}^{4})$ and $\mathbf{h}\in F$, restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}$, where $\mathbf{r}^{\prime}_{1}$ is the rank function of $U_{2,3}$, the matroid on the extreme ray of $\Gamma_{3}$. Thus, if $\mathbf{h}$ is entropic, then $a=\log{v}$ for some integer $v>0$. Now for any $\mathbf{h}\in F$ with $a=\log{v}$, let $\mathbf{h}^{\prime}$ be its restriction on $N_{3}$ and $(X_{1},X_{2},X_{3})$ be the characterizing random variable of $\mathbf{h}^{\prime}$. Let $X_{4}$ be an arbitrary random variable independent of $(X_{1},X_{2},X_{3})$ with $H(X_{4})=b$. Then it can be checked that $(X_{i},i\in N_{4})$ is the characterizing random variable of $\mathbf{h}$. Thus $\mathbf{h}$ is entropic.
For $F$ be any of the other $6$ faces and $\mathbf{h}\in F$, let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the matroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $N_{3}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i}=1,2$ is the restriction of $\mathbf{r}_{i}$ on $N_{3}$, and the rank function of $U_{2,3}$ and $U^{1,3}_{1,1}$, respectively. Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{1,3}_{1,1})$ and $a=\log{v}$ for positive $v\in\mathbb{Z}$ by [11, Theorem 1].
Now we prove the “if ” part. Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}\in F$ and $a=\log{v}$ for some integer $v\geq 0$. Let $\mathbf{h}^{\prime}$ be the restriction of $\mathbf{h}$ on $N_{3}$. Therefore, $\mathbf{h}^{\prime}\in(U_{2,3},U^{1,3}_{1,1})$. Note that $\mathbf{h}^{\prime}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are respectively the rank functions of $U_{2,3}$ and $U^{1,3}_{1,2}$. Let $(X^{\prime}_{1},X^{\prime}_{2},X^{\prime}_{3})$ be the characterizing random vector of $a\mathbf{r}_{1}$, that is, $X^{\prime}_{1}$ and $X^{\prime}_{2}$ are independent and both uniformly distributed on $\mathbb{I}_{v}$, and $X^{\prime}_{3}=X^{\prime}_{1}+X^{\prime}_{2}\mod v$. Let $(X^{\prime\prime}_{1},X^{\prime\prime}_{2},X^{\prime\prime}_{3})$ be characterizing random vector of $b\mathbf{r}_{2}$, that is, $X^{\prime\prime}_{1}$ is an arbitrary random variable with $H(X^{\prime\prime}_{1})=b$, and $X^{\prime\prime}_{2}$ and $X^{\prime\prime}_{3}$ are both constant. Let $X_{i}=(X^{\prime}_{i},X^{\prime\prime}_{i})$, $i\in N_{3}$. Then it can be checked that $(X_{1},X_{2},X_{3})$ is the characterizing random vector of $\mathbf{h}^{\prime}$.
Now for
  • $F=(U_{2,3}^{123},U_{1,1}^{1})$, let $X_{4}$ be a constant;
  • $F=(U_{2,3}^{123},U_{1,2}^{14})$, let $X_{4}=X^{\prime\prime}_{1}$;
  • $F=(\mathcal{W}_{2}^{14},U_{1,1}^{1})$, let $X_{4}=X^{\prime}_{1}$;
  • $F=(\mathcal{W}_{2}^{34},U_{1,1}^{1})$, let $X_{4}=X^{\prime}_{3}$;
  • $F=(\mathcal{W}_{2}^{14},U_{1,2}^{14})$, let $X_{4}=X_{1}$;
  • $F=(\mathcal{W}_{2}^{24},U_{1,2}^{14})$, let $X_{4}=(X^{\prime}_{2},X^{\prime\prime}_{1})$.
Then $(X_{i},i\in N_{4})$ is the characterizing random vector of $\mathbf{h}$, thus $\mathbf{h}$ is entropic. ∎
Theorem IV-B.3.
For $F=(\hat{U}_{2,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for integer $v>0$.
Proof.
Let $F$ be any of the $3$ faces and $\mathbf{h}\in F$. Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the minimal integer polymatroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i},i=1,2$ are the restriction of $\mathbf{r}_{i}$ on $\{2,3,4\}$.
By restriction, the above $3$ faces of $\Gamma_{4}$ are mapped to a subset of $\Gamma_{3}$ on $\{2,3,4\}$. The face $(\hat{U}_{2,5}^{1},U_{1,1}^{1})$ of $\Gamma_{4}$ is mapped to the extreme ray $U_{2,3}$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log v$ for integer $v>0$. For the faces $(\hat{U}_{2,5}^{1},U_{1,1}^{2})$ and $(\hat{U}_{2,5}^{1},U_{1,2}^{12})$, by restriction, they are both mapped to the face $(U_{2,3},U^{1}_{1,1})$ of $\Gamma_{3}$ on $\{2,3,4\}$, which implies that $a$ can only take $\log v$ for positive integer $v$ as well.
For the “if” part, as the whole ray $U_{1,1}^{i}$ and $U_{1,2}^{12}$ are entropic, by Lemma 2 and Theorem II-C.4, all $a=\log v$ are entropic. ∎
Theorem IV-B.4.
For $F=(\hat{U}_{2,5}^{1},U_{1,3}^{123})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b\geq\log\lceil 2^{a}\rceil$.
Proof.
Let $\mathbf{h}=a\mathbf{r}_{1}+b\mathbf{r}_{2}$, where $\mathbf{r}_{1}$ and $\mathbf{r}_{2}$ are the rank functions of the minimal integer polymatroids on the two extreme rays of the face, respectively. Restricting $\mathbf{h}$ on $\{2,3,4\}$, we obtain $\mathbf{h}^{\prime}=a\mathbf{r}^{\prime}_{1}+b\mathbf{r}^{\prime}_{2}$, where $\mathbf{r}^{\prime}_{i},i=1,2$ are the restriction of $\mathbf{r}_{i}$ on $\{2,3,4\}$. By restriction, The face $(\hat{U}_{2,5}^{1},U_{1,3}^{123})$ of $\Gamma_{4}$ is mapped to the face $(U_{2,3},U^{12}_{1,2})$ of $\Gamma_{3}$ on $\{2,3,4\}$, Thus, $\mathbf{h}^{\prime}\in(U_{2,3},U^{12}_{1,2})$ and $a+b\geq\log\lceil 2^{a}\rceil$ by [10, Theorem 1].
For the “if ” part , we only need to show that $\mathbf{h}$ with $a+b=\log v$ is entropic for each positive integer $v$, then the theorem is implied by Lemma 2 and the fact that $b\mathbf{r}_{2}$ is entropic for all $b\geq 0$. Let $X_{2}$ be uniformly distributed on $\mathbb{I}_{v}$. Let $X_{2}$ and $X_{4}$ be independent, and $X_{4}$ be distributed on $\mathbb{I}_{v}$ such that $H(X_{4})=a$. Let $X_{1}=vX_{2}+X_{4}$, $X_{3}=X_{2}+X_{4}\mod v$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,\log v-a)$. ∎
Theorem IV-B.5.
For $F=(\hat{U}_{3,5}^{1},U_{1,1}^{1})$, $(\hat{U}_{3,5}^{1},U_{1,1}^{2})$, or $(\hat{U}_{3,5}^{1},U_{1,2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for some integer $v>0$.
Proof.
If $\mathbf{h}^{\prime}$ is entropic, where $\mathbf{h}^{\prime}$ is obtained by restricting $\mathbf{h}=(a,b)$ on $\{1,3,4\}$, its characterizing random vector $(X_{i},i\in\{1,3,4\})$ satisfies
$H(X_{134})$ $=H(X_{13})=H(X_{14})$
and
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),\text{ distinct }i,j\in\{1,3,4\}$
For $x_{134}$ with $p(x_{134})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})=p(x_{1},x_{4})$
$=p(x_{1})p(x_{3})=p(x_{1})p(x_{4}),$
which implies $p(x_{3})=p(x_{4})$. Since $X_{3}$ and $X_{4}$ are independent, by Lemma 1, they are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and $H(X_{3})=H(X_{4})=\log v$, where $v=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$. As $(X_{i},i\in\{1,3,4\})$ is its characterizing random vector, we have
$H(X_{3})$ $=H(X_{4})=a.$
Thus $a$ can only take $\log v$.
For the “if” part, it can be proven by Lemma 2 and the fact that $a=\log{v}$ in $\hat{U}_{3,5}^{1}$ and whole ray $U^{1}_{1,1}$, $U^{2}_{1,1}$ and $U_{1,2}^{12}$ are entropic. ∎
IV-C Entropy functions on the faces involving $U_{2,3}^{123}$
In this subsection, we characterize entropy functions on the $2$-dimensional faces of $\Gamma_{4}$ with one of its extreme rays $U_{2,3}^{123}$ except those characterized in Subsection IV-B. As introduced in [7], uniform matroid $U_{2,3}$ corresponding to Latin squares or orthogonal arrays $\mathrm{OA}(2,3)$, the characterizing random vectors of these entropy functions are distributed on the combinatorial structures extending Latin squares.
For $\mathbf{h}\in(U_{2,3}^{123},U_{1,3}^{124})$, if we restrict it on $N_{3}$, we will see that its restriction $\mathbf{h}^{\prime}\in(U_{2,3},U^{12,3}_{1,2})$. However, we will prove in Theorem IV-C.1 that the shape of the entropy functions on $(U_{2,3}^{123},U_{1,3}^{124})$ is the same as $(U_{2,3},U_{1,1}^{1,3})$, which is depicted in Figure 3. For $(U_{2,3}^{123},U_{2,3}^{124})$ and $(U_{2,3}^{123},U_{1,4})$, they are characterized in Theorems IV-C.2 and IV-C.3, and depicted in Figure 4 and 5, respectively.
Theorem IV-C.1.
For $F=(U_{2,3}^{123},U_{1,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$.
Figure 3: The face $(U_{2,3}^{123},U_{1,3}^{124})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{3j})$ $=H(X_{3})+H(X_{j}),j\in N_{4}\setminus\{3\},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),k\neq 3,$
$\qquad\qquad\qquad\qquad\{i,j,k\}\neq\{1,2,3\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,2,3\}$
$\qquad\qquad\quad|K|=2,\{i,j\}=N_{4}\setminus K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.C.1)
$=p(x_{1},x_{2},x_{4})$ (IV.C.2)
$=p(x_{1},x_{3},x_{4})$ (IV.C.3)
$=p(x_{2},x_{3},x_{4}),$ (IV.C.4)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.C.5)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.C.6)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.C.7)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (IV.C.8)
$p(x_{1},x_{2})p(x_{2},x_{4})$ $=p(x_{2})p(x_{1},x_{2},x_{4}),$ (IV.C.9)
$p(x_{1},x_{4})p(x_{2},x_{4})$ $=p(x_{4})p(x_{1},x_{2},x_{4}),$ (IV.C.10)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (IV.C.11)
$p(x_{1},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{1},x_{3},x_{4}),$ (IV.C.12)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.C.13)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (IV.C.14)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (IV.C.15)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (IV.C.16)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}).$ (IV.C.17)
By (IV.C.1)-(IV.C.4) and (IV.C.15)-(IV.C.17), we have
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.C.18)
$=p(x_{1},x_{2})$ (IV.C.19)
$=p(x_{1},x_{3})$ (IV.C.20)
$=p(x_{2},x_{3}),$ (IV.C.21)
which implies that each of the random variables $X_{i},i\in N_{3}$ is a function of the other two.
Consider the bipartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}$ and $(x_{1},x_{2})\in E$ if and only if $p(x_{1},x_{2})>0$. By (IV.C.19), for each $(x_{1},x_{2})\in E$, we color it by the unique $x_{3}\in\mathcal{X}_{3}$. Assume $|\mathcal{X}_{3}|=v$ and $G$ has $t$ connected components. We denote the number of the vertices of $\mathcal{X}_{i}$ in the connected component $C_{j}$ by $n_{i}^{(j)},i=1,2,j=1,2,\cdots,t$ and the probability mass of $C_{j}$ by $p_{j}$, that is, the probability of the event that the random vector takes a tuple in $C_{j}$. In light of (IV.C.20) and (IV.C.21), equating (IV.C.5) and (IV.C.6), we have
$p(x_{1})=p(x_{2}).$ (IV.C.22)
Let $(x^{\prime}_{1},x_{2})\in E$ be another edge adjacent to $x_{2}$, by the same argument, we have
$p(x_{1}^{\prime})=p(x_{2}).$ (IV.C.23)
As $x^{\prime}_{1}$ is arbitrarily chosen, all vertices adjacent to $x_{2}$ have the same probability mass. Then by symmetry, so do all vertices in the connected component $C_{j}$, and $n_{1}^{(j)}=n_{2}^{(j)}$ for each $j$. Herein, we simplify $n_{1}^{(j)}$ to $n^{(j)}$. By (IV.C.6), $X_{2}$ and $X_{3}$ are independent, which implies that all colors $x_{3}\in\mathcal{X}_{3}$ appear in the edges incident to a vertex $x_{2}\in\mathcal{X}_{2}$. Hence,
$n^{(j)}\geq v$ (IV.C.24)
for each $j$. Then,
$H(X_{1})$ $=-\sum_{i=1}^{t}n^{(i)}\frac{p_{i}}{n^{(i)}}\log(\frac{p_{i}}{n^{(i)}})$ (IV.C.25)
$=H(p_{1},p_{2},\ldots,p_{t})+\sum_{i=1}^{t}p_{i}\log{n^{(i)}}.$ (IV.C.26)
By (IV.C.18), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (IV.C.15), we obtain
$p(x_{1},x_{2})=p(x_{1},x_{2},x_{4}).$ (IV.C.27)
Together with (IV.C.8) and (IV.C.9), we have
$p(x_{1})$ $=p(x_{1},x_{4})$ (IV.C.28)
and
$p(x_{2})$ $=p(x_{2},x_{4}).$ (IV.C.29)
By (IV.C.27), $X_{4}$ is a function of $X_{1}$ and $X_{2}$, and so we recolor each $(x_{1},x_{2})\in E$ by $x_{4}\in\mathcal{X}_{4}$. By (IV.C.28) and (IV.C.29), all the edges adjacent to a vertex have the same color, which implies edges in the whole component have the same color. Hence
$H(X_{4})\leq H(p_{1},p_{2},\ldots,p_{t}).$ (IV.C.30)
The inequality holds strictly when there exist two components of $G$ with the same color.
As $\mathbf{h}\in(U_{2,3}^{123},U_{1,3}^{124})$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=a+b,$ (IV.C.31)
$H(X_{2})$ $=a+b,$ (IV.C.32)
$H(X_{3})$ $=a,$ (IV.C.33)
$H(X_{4})$ $=b.$ (IV.C.34)
Replacing the right hand side of (IV.C.31) into (IV.C.26) yields
$a+b=H(p_{1},p_{2},\ldots,p_{t})+\sum\limits_{i=1}^{t}p_{i}\log{n}^{(i)}.$ (IV.C.35)
By (IV.C.30) and (IV.C.34), canceling $b$ and $H(p_{1},p_{2},\ldots,p_{t})$ at either side of (IV.C.35), we obtain
$a\geq\sum\limits_{i=1}^{t}p_{i}\log{n^{(i)}},$ (IV.C.36)
which together with (IV.C.24) yields
$a\geq\sum\limits_{i=1}^{t}p_{i}\log{v}=\log{v}.$ (IV.C.37)
Note that $H(X_{3})\leq\log{v}$. Thus, by (IV.C.33) and (IV.C.37),
$\log{v}\geq a\geq\log{v},$ (IV.C.38)
which implies that $a$ can only take $\log{v}$ for some positive integer $k$.
The “if” part of the theorem is immediately implied by Lemma 2 and the fact that the $a=\log{v}$ in $U_{2,3}^{123}$ and whole ray $U_{1,3}^{124}$ are entropic. ∎
Theorem IV-C.2.
For $F=(U_{2,3}^{123},U_{2,3}^{124})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{v_{1}}$ and $b=\log{v_{2}}$ for positive integers $v_{1},v_{2}$.
Figure 4: The face $(U_{2,3}^{123},U_{2,3}^{124})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),\ i,j\in N_{4},i\neq j,$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),$
$\qquad\qquad\{i,j,k\}\neq{\{1,2,3\},\{1,2,4\}},$
$H(X_{123})+H(X_{124})$ $=H(X_{12})+H(X_{N_{4}}).$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.C.39)
$=p(x_{1},x_{2},x_{4})$ (IV.C.40)
$=p(x_{1},x_{3},x_{4})$ (IV.C.41)
$=p(x_{2},x_{3},x_{4}),$ (IV.C.42)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.C.43)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.C.44)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.C.45)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.C.46)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.C.47)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.C.48)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (IV.C.49)
$p(x_{1},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{1},x_{3},x_{4}),$ (IV.C.50)
$p(x_{1},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{1},x_{3},x_{4}),$ (IV.C.51)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.C.52)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.C.53)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (IV.C.54)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (IV.C.55)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}).$ (IV.C.56)
By (IV.C.40), canceling $p(x_{1},x_{2},x_{4})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (IV.C.56), we have
$p(x_{1},x_{2},x_{3})=p(x_{1},x_{2}).$ (IV.C.57)
Equating (IV.C.43) and (IV.C.57) yields
$p(x_{1},x_{2},x_{3})=p(x_{1})p(x_{2}).$ (IV.C.58)
By (IV.C.44) and (IV.C.45), replacing $p(x_{1},x_{3})$ and $p(x_{1},x_{4})$ by $p(x_{1})p(x_{3})$ and $p(x_{1})p(x_{4})$ in (IV.C.49), we have
$p(x_{1},x_{3},x_{4})=p(x_{1})p(x_{3})p(x_{4}).$ (IV.C.59)
By the same argument, we have
$p(x_{2},x_{3},x_{4})=p(x_{2})p(x_{3})p(x_{4}).$ (IV.C.60)
By (IV.C.39) and (IV.C.42), equating (IV.C.58) and (IV.C.60) yields
$p(x_{1})=p(x_{3})p(x_{4}).$ (IV.C.61)
By (IV.C.59), $X_{1}$, $X_{3}$ and $X_{4}$ are mutually independent. Let $x^{\prime}_{3}\in\mathcal{X}_{3}$ with $x^{\prime}_{3}\neq x_{3}$, then
$p(x_{1},x^{\prime}_{3},x_{4})=p(x_{1})p(x^{\prime}_{3})p(x_{4})>0.$ (IV.C.62)
As
$p(x_{1},x^{\prime}_{3},x_{4})=\sum_{x_{2}}p(x_{1},x_{2},x^{\prime}_{3},x_{4}),$ (IV.C.63)
there exists $x’_{2}\in\mathcal{X}_{2}$ such that
$p(x_{1},x^{\prime}_{2},x’_{3},x_{4})>0.$ (IV.C.64)
By the same argument above, we obtain
$p(x_{1})=p(x_{3}{{}^{\prime}})p(x_{4}).$ (IV.C.65)
Equating (IV.C.61) and (IV.C.65), we have
$p(x_{3})=p(x_{3}{{}^{\prime}}).$ (IV.C.66)
As (IV.C.66) holds for any $x_{3}^{\prime}\in\mathcal{X}_{3}$, $X_{3}$ is uniformly distributed on $\mathcal{X}_{3}$, and so $H(X_{3})=\log{v_{1}}$ where $v_{1}=|\mathcal{X}_{3}|$. By the same argument, $X_{4}$ is uniform, and so $H(X_{4})=\log{v_{2}}$ where $v_{2}=|\mathcal{X}_{4}|$.
For the “if” part of the theorem, as $\mathbf{h}\in U_{2,3}^{123}$ and $\mathbf{h^{\prime}}\in U_{2,3}^{124}$ are entropic if $a=\log{v_{1}},b=\log{v_{2}}$ for integers $v_{1},v_{2}>0$ [35]. By Lemma 2, all $(\log{v_{1}},\log{v_{2}})$ are entropic. ∎
Theorem IV-C.3.
For $F=(U_{2,3}^{123},U_{1,4})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a\geq 0,b>0$ or $(a,b)=(\log{v},0)$ for positive integer $v$.
[…]
Figure 5: The face $(U_{2,3}^{123},U_{1,4})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),\{i,j,k\}\neq{\{1,2,3\}},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,2,3\},$
$\qquad\qquad\quad|K|=2,\{i,j\}=N_{4}\setminus K.$
For $\left(x_{1},x_{2},x_{3},x_{4}\right)\in\mathcal{X}_{N_{4}}$ with $p(x_{1},x_{2},x_{3},x_{4})>0$, the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.C.67)
$=p(x_{1},x_{2},x_{4})$ (IV.C.68)
$=p(x_{1},x_{3},x_{4})$ (IV.C.69)
$=p(x_{2},x_{3},x_{4}),$ (IV.C.70)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (IV.C.71)
$p(x_{1},x_{2})p(x_{2},x_{4})$ $=p(x_{2})p(x_{1},x_{2},x_{4}),$ (IV.C.72)
$p(x_{1},x_{4})p(x_{2},x_{4})$ $=p(x_{4})p(x_{1},x_{2},x_{4}),$ (IV.C.73)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (IV.C.74)
$p(x_{1},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{1},x_{3},x_{4}),$ (IV.C.75)
$p(x_{1},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{1},x_{3},x_{4}),$ (IV.C.76)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.C.77)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.C.78)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (IV.C.79)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (IV.C.80)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (IV.C.81)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}).$ (IV.C.82)
By (IV.C.67), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (IV.C.80), we obtain
$p(x_{1},x_{3},x_{4})=p(x_{1},x_{2}).$ (IV.C.83)
Equating (IV.C.69) and (IV.C.83) yields,
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2}).$ (IV.C.84)
By the same argument,
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{3})$ (IV.C.85)
$=p(x_{2},x_{3}).$ (IV.C.86)
In light of (IV.C.68) and (IV.C.84), we have
$p(x_{1},x_{2})=p(x_{1},x_{2},x_{4}).$ (IV.C.87)
By (IV.C.87), canceling $p(x_{1},x_{2})$ and $p(x_{1},x_{2},x_{4})$ on either side of (IV.C.71), we have
$p(x_{1},x_{4})=p(x_{1}).$ (IV.C.88)
Similarly,
$p(x_{2},x_{4})$ $=p(x_{2}),$ (IV.C.89)
$p(x_{3},x_{4})$ $=p(x_{3}).$ (IV.C.90)
By (IV.C.88)-(IV.C.90), replacing $p(x_{1},x_{4})$ by $p(x_{1})$, $p(x_{2},x_{4})$ by $p(x_{2})$ and $p(x_{3},x_{4})$ by $p(x_{3})$ in (IV.C.73), (IV.C.76) and (IV.C.79), we obtain
$p(x_{1})p(x_{2})=p(x_{4})p(x_{1},x_{2},x_{4}),$ (IV.C.91)
$p(x_{1})p(x_{3})=p(x_{4})p(x_{1},x_{3},x_{4}),$ (IV.C.92)
$p(x_{2})p(x_{3})=p(x_{4})p(x_{2},x_{3},x_{4}).$ (IV.C.93)
Then, by (IV.C.68)-(IV.C.70), (IV.C.91)-(IV.C.93) we conclude
$p(x_{1})=p(x_{2})=p(x_{3}).$ (IV.C.94)
Consider the tripartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}$ and $(x_{i},x_{j})\in E$ if and only if $p(x_{i},x_{j})>0$, ${i,j}\in N_{3},i\neq j$. By (IV.C.84)-(IV.C.86), for each $(x_{i},x_{j})\in E$, we color it by the unique $x_{4}\in\mathcal{X}_{4}$. Assume $G$ has $t$ connected components. We denote the number of the vertices of $\mathcal{X}_{i}$ in the connected component $C_{j}$ by $n_{i}^{(j)},i=1,2,3,j=1,2,\cdots,t$ and the probability mass of $C_{j}$ by $p_{j}$. Now for any $(x_{1},x_{2})\in\mathcal{X}_{1}\times\mathcal{X}_{2}$ with $p(x_{1},x_{2})>0$, by (IV.C.84), there exist unique $x_{3}$ and $x_{4}$ with $p(x_{1},x_{2},x_{3},x_{4})>0$, and so by (IV.C.94), $p(x_{1})=p(x_{2})$. Similarly, for $p(x_{1},x_{3})>0$ and $p(x_{2},x_{3})>0$, we have $p(x_{1})=p(x_{3})$ and $p(x_{2})=p(x_{3})$. Then we conclude that $p(x_{i})$$p(x_{j})$ when $p(x_{i},x_{j})>0$. Thus each vertex in a component of $G$ has the same probability mass. By (IV.C.88)-(IV.C.90), all the edges adjacent to a vertex have the same color, which implies edges in the whole component have the same color. Hence
$p(x_{4})\geq p_{j}.$ (IV.C.95)
The inequality strictly holds when there exist two components sharing the same color.
By (IV.C.68), replacing $p(x_{1},x_{2},x_{4})$ by $p(x_{1},x_{2},x_{3},x_{4})$ in (IV.C.91), we obtain
$p(x_{1})p(x_{2})=p(x_{4})p(x_{1},x_{2},x_{3},x_{4}).$ (IV.C.96)
By (IV.C.96), it can be seen that $p(x_{1},x_{2},x_{3},x_{4})$ are all equal for $(x_{1},x_{2},x_{3},x_{4})\in\mathcal{X}^{(j)}_{N_{4}}$, where $\mathcal{X}^{(j)}_{N_{4}}$ is the alphabet $\mathcal{X}_{N_{4}}$ in the connected component $C_{j}$. Summing up over all $p(x_{1})$ in $C_{j}$,
$p_{j}=\sum\limits_{x_{1}\in\mathcal{X}^{(j)}_{1}}p(x_{1}).$ (IV.C.97)
Since each vertex has the same probability mass in $C_{j}$, the number of the vertices $n_{1}^{(j)}=n_{2}^{(j)}=n_{3}^{(j)}=\dfrac{p_{j}}{p(x_{1})}$ where $p_{j}$ is the probability mass of $C_{j}$. We simplify $n_{1}^{(j)}$ to $n^{(j)}$, then $p(x_{1})=p(x_{2})=\dfrac{p_{j}}{n^{(j)}}$. Then substituting $p(x_{1})$ and $p(x_{2})$ by $\dfrac{p_{j}}{n^{(j)}}$ in (IV.C.96), we have
$p(x_{1},x_{2},x_{3},x_{4})=\dfrac{p_{j}^{2}}{({n^{(j)})}^{2}p(x_{4})}.$ (IV.C.98)
Summing up all probability mass in $C_{j}$,
$p_{j}=\sum\limits_{(x_{1},x_{2},x_{3},x_{4})\in\mathcal{X}^{(j)}_{N_{4}}:\atop p(x_{1},x_{2},x_{3},x_{4})>0}p(x_{1},x_{2},x_{3},x_{4}).$ (IV.C.99)
As $p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})$,
$p_{j}=\sum\limits_{(x_{1},x_{2})\in\mathcal{X}^{(j)}_{N_{2}}:\atop p(x_{1},x_{2})>0}p(x_{1},x_{2}).$ (IV.C.100)
Since $n_{1}^{(j)}=n_{2}^{(j)}=n^{(j)}$, we have
$|\{x_{12}\in\mathcal{X}^{(j)}_{N_{2}}:p(x_{12})>0\}|\leq|\mathcal{X}^{(j)}_{N_{2}}|=({n^{(j)})}^{2}.$ (IV.C.101)
Since $p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})$, $p(x_{1},x_{2})$ are equal for all $(x_{1},x_{2})\in\mathcal{X}^{(j)}_{N_{2}}$. Then by (IV.C.100) and (IV.C.101),
$p_{j}\leq({n^{(j)})}^{2}p(x_{1},x_{2}).$ (IV.C.102)
By (IV.C.84) , (IV.C.98) and (IV.C.102), we have
$p_{j}\leq({n^{(j)})}^{2}p(x_{1},x_{2})=\dfrac{p_{j}^{2}}{p(x_{4})},$ (IV.C.103)
which implies $p_{j}\geq p(x_{4}).$ Together with (IV.C.95), we have
$p_{j}=p(x_{4}).$ (IV.C.104)
Therefore, there exist no two components sharing the same color, that is
$t=|\mathcal{X}_{4}|.$ (IV.C.105)
By (IV.C.104), (IV.C.103) holds with equality which implies (IV.C.102) and (IV.C.101) hold with equality as well. Hence for any $x_{1}\in\mathcal{X}^{(j)}_{1}$, $x_{2}\in\mathcal{X}^{(j)}_{2}$, $x_{1}$ is adjacent to $x_{2}$. By the same argument, we can conclude that $x_{i}\in\mathcal{X}^{(j)}_{i}$ is adjacent to the vertex in $\mathcal{X}^{(j)}_{N_{3}-i}$ for $i=1,2,3$.
Now we prove that for all $\mathbf{h}=(a,b)\in F$, $a\geq 0,b>0$ is entropic. For $a=\log v$ with some positive $v$ and $b>0$, the fact that $(a,b)$ is entropic can be implied by Lemma2. Here we give a constructive proof of the case. Let $X_{4}$ be any random variable on $\mathcal{X}_{4}=\{0,1,\ldots,t-1\}$ with $H(X_{4})=b$. Note that its probability mass is $p_{i},i\in\mathcal{X}_{4}$. Thus
$b=\sum_{i=}^{t-1}-p_{i}\log p_{i}.$ (IV.C.106)
Now assume in each connected component $C_{j}$ of $G$, $\mathcal{X}_{i}^{(j)}$ has $k$ element, that is, $n^{(j)}=k$ for all $j\in\mathcal{X}_{4}$. Let $X^{(j)}_{1}$ and $X^{(j)}_{2}$ be independent and uniformly distributed on $\mathbb{I}_{v}$, and $X^{(j)}_{3}=X^{(j)}_{1}+X^{(j)}_{2}\mod v$. Let $X_{i}=(X^{(j)}_{i},j\in\mathcal{X}_{4})$, $i=1,2,3$. Then
$H(X_{1})$ $=H(X_{2})=H(X_{3})=\sum_{j=0}^{t-1}-p_{j}\log\frac{p_{j}}{k}$ (IV.C.107)
$=\sum\limits_{j=0}^{t-1}p_{j}\log{v}+\sum_{j=0}^{t-1}-p_{j}\log p_{j}$ (IV.C.108)
$=\log{v}+H(X_{4})=a+b.$ (IV.C.109)
It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,b)$.
Consider $a\neq\log{v},b>0$. Let $c=\lfloor 2^{a}\rfloor$. Let $X_{4}$ be a random variable on $\mathcal{X}_{4}=\{0,1,\ldots,t-1\}$ with $H(X_{4})=b$. Let $n^{(j)}=c$ for $j\in\mathcal{X}_{4}\setminus\{t-1\}$ and $n^{(t-1)}=c+l$ for some sufficiently large $l$. Let $X^{(j)}_{1}$ and $X^{(j)}_{2}$ be independent and uniformly distributed on $\{0,1,\dots,c-1\}$ for $j\in\mathcal{X}_{4}\setminus\{t-1\}$ and $X^{(j)}_{3}=X^{(j)}_{1}+X^{(j)}_{2}\mod c$. Let $X^{(t-1)}_{1}$ and $X^{(t-1)}_{2}$ be independent and uniformly distributed on $\{0,1,\dots,c+l-1\}$ and $X^{(t-1)}_{3}=X^{(t-1)}_{1}+X^{(t-1)}_{2}\mod c+l$. Let $X_{i}=(X^{(j)}_{i},j\in\mathcal{X}_{4})$, $i=1,2,3$. Then,
$H(X_{1})$ $=H(X_{2})=H(X_{3})$ (IV.C.110)
$=\sum\limits_{i=0}^{t-2}(-p_{i}\log(\dfrac{p_{i}}{c}))-p_{t-1}\log\dfrac{p_{t-1}}{c+l}$ (IV.C.111)
$=\sum\limits_{i=0}^{t-1}-p_{i}\log(p_{i})+\sum\limits_{i=0}^{t-2}p_{i}\log c+p_{t-1}\log(c+l)$ (IV.C.112)
$=H(X_{4})+\log{c}({1-p_{t-1}})+\log{(c+l)}p_{t-1}.$ (IV.C.113)
Note that $l$ is sufficiently large, so let
$p_{t}=\dfrac{a-\log{c}}{\log{(c+l)}-\log{c}},$ (IV.C.114)
we have
$H(X_{1})$ $=H(X_{2})=H(X_{3})=H(X_{4})+a.$ (IV.C.115)
It can be checked that the entropic function of such constructed $(X_{i},i\in N_{4})$ is $(a,b)$.
Hence, all $\mathbf{h}=(a,b)\in F$ are entropic when $a\geq 0,b>0$. Note that $\mathbf{h}\in U_{2,3}^{123}$ are entropic if and only if $a=\log{v}$ for positive integer $v$. The theorem is proved. ∎
Theorem IV-C.3 can be considered as a corollary of [32, Lemma 4], which is restated as [36, Lemma 6].
IV-D Entropy functions on the faces involving partitions
In this subsection, we characterize the faces where the distribution of the entropy function’s characterizing random vector involves number partitions. A $t$-partition of a positive $v$ with $0\leq t\leq v-1$ is a $t$-tuple $\bm{\alpha}=(\alpha_{0},\alpha_{1},\dots,\alpha_{t-1})$ with $\alpha_{i}$ positive integer and $\sum_{i=0}^{t-1}\alpha_{i}=v$. The family of all such partitions of $v$ is denoted by $\mathcal{P}(v)$ and $H(\bm{\alpha})$ reprensent the shannon entropy $H(\frac{\alpha_{i}}{v}:i\in\mathbb{I}_{t})$.
Theorem IV-D.1.
For $F=(\mathcal{W}_{2}^{12},U_{2,3}^{134})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log{v},a=H(\dfrac{\alpha_{1}}{v},\dfrac{\alpha_{2}}{v},\dots,\dfrac{\alpha_{t}}{v})$ , where integer $v>0$ and $(\alpha_{0},\alpha_{1},\dots,\alpha_{t-1})$ is a partition of $v$.
[…]
Figure 6: The face $(\mathcal{W}_{2}^{12},U_{2,3}^{134})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),\{i,j\}\neq\{1,2\}$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{i,j,k}),k\in\{1,2\}\subseteq{\{i,j,k\}},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),K\subseteq\{1,3,4\},$
$\qquad\qquad\qquad|K|=2,\{i,j\}=N_{4}\setminus K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.D.1)
$=p(x_{1},x_{2},x_{4})$ (IV.D.2)
$=p(x_{1},x_{3},x_{4})$ (IV.D.3)
$=p(x_{2},x_{3},x_{4}),$ (IV.D.4)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.D.5)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.D.6)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.D.7)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.D.8)
$p(x_{3},x_{4})$ $=p(x_{4})p(x_{4}),$ (IV.D.9)
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (IV.D.10)
$p(x_{1},x_{2})p(x_{2},x_{3})$ $=p(x_{2})p(x_{1},x_{2},x_{3}),$ (IV.D.11)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (IV.D.12)
$p(x_{1},x_{2})p(x_{2},x_{4})$ $=p(x_{2})p(x_{1},x_{2},x_{4}),$ (IV.D.13)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (IV.D.14)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}),$ (IV.D.15)
$p(x_{2},x_{3},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (IV.D.16)
By (IV.D.1), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (IV.D.14) yields
$p(x_{1},x_{3},x_{4})=p(x_{1},x_{3}).$ (IV.D.17)
Together with (IV.D.5), we have
$p(x_{1},x_{3},x_{4})=p(x_{1})(x_{3}).$ (IV.D.18)
By the same argument,
$p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})=p(x_{1})p(x_{4})$ (IV.D.19)
$=p(x_{3},x_{4})=p(x_{3})p(x_{4}).$ (IV.D.20)
Equating (IV.D.18), (IV.D.19) and (IV.D.20), we have
$p(x_{1})=p(x_{3})=p(x_{4}).$ (IV.D.21)
By (IV.D.17)-(IV.D.20), $X_{1},X_{3}$ and $X_{4}$ are pairwise independent and each one is a function of the other two. Thus, $(X_{1},X_{3},X_{4})$ is the characterizing random vector of an entropy function on $U_{2,3}$, an extreme ray of $\Gamma_{3}$, and $X_{i}$ is uniformly distributed on $\mathcal{X}_{i}$ and $H(X_{i})=\log v$ where $|\mathcal{X}_{i}|=v$ for $i=1,3,4$.
As $\mathbf{h}\in F$, $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H\left(X_{1}\right)$ $=a+b=\log{v},$ (IV.D.22)
$H(X_{2})$ $=a,$ (IV.D.23)
$H\left(X_{3}\right)$ $=a+b=\log{v},$ (IV.D.24)
$H\left(X_{4}\right)$ $=a+b=\log{v}.$ (IV.D.25)
By (IV.D.1), (IV.D.3) and (IV.D.17),
$p(x_{1},x_{2},x_{3})=p(x_{1},x_{3}).$ (IV.D.26)
By (IV.D.26), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{3})$ on either side of (IV.D.10), we have
$p(x_{1})=p(x_{1},x_{2}).$ (IV.D.27)
By (IV.D.27), $X_{2}$ is a function of $X_{1}$. Thus there exists a unique $x_{2}\in\mathcal{X}_{2}$ with $p(x_{1},x_{2})>0$, then $\mid\mathcal{X}_{2}\mid\leq\mid\mathcal{X}_{1}\mid=v$ and
$p(x_{2})=\sum\limits_{x_{1}}p(x_{1},x_{2})=\dfrac{\alpha\left(x_{2}\right)}{v},$ (IV.D.28)
[…]
Now we prove all these points are entropic. Let $X_{1},X_{3}$ be indenpendent and uniformly distributed on $\mathcal{X}_{1}=\mathcal{X}_{3}=\mathbb{I}_{v}$. Let $X_{4}=X_{1}+X_{3}$ mod $v$. It can be checked that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}=\mathbb{I}_{v}$. Let $\left(A(x_{2}):x_{2}\in\mathcal{X}_{2}\right)$ be a partition of $\mathbb{I}_{v}$, where $|A(x_{2})|=\alpha(x_{2}).$ Let $X_{2}$ be a function $\varphi$ of $X_{1}$, where $\varphi(x_{1})=x_{2}$ if $x_{1}\in A(x_{2})$. Then it can be checked that $(X_{i},i\in N_{4})$ is the characterizing random vector of $\mathbf{h}\in F$. ∎
IV-E Entropy functions on the faces involving $U_{3,4}$
In this subsection, we characterize entropy functions on the 55 22-dimensional faces of $\Gamma_{4}$ with one of its extreme rays $U_{3,4}$. As introduced in [7], uniform matroid $U_{3,4}$ corresponding to Latin cubes or orthogonal arrays $\mathrm{OA}(3,4)$, the characterizing random vectors of these entropy functions are distributed on the combinatorial structures extending Latin cubes.
Theorem IV-E.1.
For $F=(U_{3,4},U^{1}_{1,1})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}\setminus\{1\},$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),\ i,j\in N_{4},i<j,$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),\{i,j,k\}\subseteq N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.E.1)
$=p(x_{1},x_{2},x_{4})$ (IV.E.2)
$=p(x_{1},x_{3},x_{4}),$ (IV.E.3)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.E.4)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.E.5)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.E.6)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.E.7)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.E.8)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.E.9)
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (IV.E.10)
$p(x_{1},x_{2})p(x_{2},x_{3})$ $=p(x_{2})p(x_{1},x_{2},x_{3}),$ (IV.E.11)
$p(x_{1},x_{3})p(x_{2},x_{3})$ $=p(x_{3})p(x_{1},x_{2},x_{3}),$ (IV.E.12)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (IV.E.13)
$p(x_{1},x_{2})p(x_{2},x_{4})$ $=p(x_{2})p(x_{1},x_{2},x_{4}),$ (IV.E.14)
$p(x_{1},x_{4})p(x_{2},x_{4})$ $=p(x_{4})p(x_{1},x_{2},x_{4}),$ (IV.E.15)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (IV.E.16)
$p(x_{1},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{1},x_{3},x_{4}),$ (IV.E.17)
$p(x_{1},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{1},x_{3},x_{4}),$ (IV.E.18)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.E.19)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.E.20)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}).$ (IV.E.21)
Since $p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})$, together with (IV.E.10) and (IV.E.13), we have
$p(x_{1},x_{3})=p(x_{1},x_{4}).$ (IV.E.22)
In light of (IV.E.5) and (IV.E.6), replacing $p(x_{1},x_{3})$ and $p(x_{1},x_{4})$ by $p(x_{1})p(x_{3})$ and $p(x_{1})p(x_{4})$ in (IV.E.22), we obtain
$p(x_{3})=p(x_{4}).$ (IV.E.23)
Note that $X_{3}$ and $X_{4}$ are independent, by Lemma 1, $X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and so $H(X_{3})=H(X_{4})=\log v$ where $v=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$.
For the “if” part, it can be proved by Lemma 2 and the fact that $a=\log{v}$ on the ray $U_{3,4}$, and whole ray $U^{1}_{1,1}$ are entropic. ∎
Theorem IV-E.2.
For $F=(U_{3,4},U^{12}_{1,2})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i<j,\{i,j\}\neq\{1,2\}$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),\{i,j\}\neq\{1,2\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.E.24)
$=p(x_{1},x_{2},x_{4})$ (IV.E.25)
$=p(x_{1},x_{3},x_{4})$ (IV.E.26)
$=p(x_{2},x_{3},x_{4}),$ (IV.E.27)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.E.28)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.E.29)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.E.30)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.E.31)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.E.32)
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (IV.E.33)
$p(x_{1},x_{2})p(x_{2},x_{3})$ $=p(x_{2})p(x_{1},x_{2},x_{3}),$ (IV.E.34)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (IV.E.35)
$p(x_{1},x_{2})p(x_{2},x_{4})$ $=p(x_{2})p(x_{1},x_{2},x_{4}),$ (IV.E.36)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (IV.E.37)
$p(x_{1},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{1},x_{3},x_{4}),$ (IV.E.38)
$p(x_{1},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{1},x_{3},x_{4}),$ (IV.E.39)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.E.40)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.E.41)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}).$ (IV.E.42)
Since $p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})$, together with (IV.E.33) and (IV.E.35), we have
$p(x_{1},x_{3})=p(x_{1},x_{4}).$ (IV.E.43)
By (IV.E.28) and (IV.E.29), replacing $p(x_{1},x_{3})$ and $p(x_{1},x_{4})$ by $p(x_{1})p(x_{3})$ and $p(x_{1})p(x_{4})$ in (IV.E.43),
$p(x_{3})=p(x_{4}).$ (IV.E.44)
By (IV.E.32), $X_{3}$ and $X_{4}$ are independent. Due to Lemma 1, we conclude that $X_{3}$ and $X_{4}$ are uniformly distributed on $\mathcal{X}_{3}$ and $\mathcal{X}_{4}$, respectively, and so $H(X_{3})=H(X_{4})=\log v$ where $v=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$.
On the other hand, the “if” part of the theorem can be proved by Lemma 2 and the fact that $a=\log{v}$ on the ray $U_{3,4}$ and whole ray $U^{12}_{1,2}$ are entropic. ∎
Theorem IV-E.3.
For $F=(U_{3,4},U^{123}_{2,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if a+ba+b= $\log{v}$ for integer $v>0$ .
Figure 7: The face $(U_{3,4},U^{123}_{2,3})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4}$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i<j,i,j\in N_{4},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),\{i,j,k\}\neq\{1,2,3\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.E.45)
$=p(x_{1},x_{2},x_{4})$ (IV.E.46)
$=p(x_{1},x_{3},x_{4})$ (IV.E.47)
$=p(x_{2},x_{3},x_{4}),$ (IV.E.48)
$p(x_{1}.x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.E.49)
$p(x_{1}.x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.E.50)
$p(x_{1}.x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.E.51)
$p(x_{2}.x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.E.52)
$p(x_{2}.x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.E.53)
$p(x_{3}.x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.E.54)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (IV.E.55)
$p(x_{1},x_{2})p(x_{2},x_{4})$ $=p(x_{2})p(x_{1},x_{2},x_{4}),$ (IV.E.56)
$p(x_{1},x_{4})p(x_{2},x_{4})$ $=p(x_{4})p(x_{1},x_{2},x_{4}),$ (IV.E.57)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (IV.E.58)
$p(x_{1},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{1},x_{3},x_{4}),$ (IV.E.59)
$p(x_{1},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{1},x_{3},x_{4}),$ (IV.E.60)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.E.61)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.E.62)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}).$ (IV.E.63)
By (IV.E.49) and (IV.E.51), replacing $p(x_{1},x_{2})$ and $p(x_{1},x_{4})$ by $p(x_{1})p(x_{2})$ and $p(x_{1})p(x_{4})$ in (IV.E.55), we have
$p(x_{1},x_{2},x_{4})=p(x_{1})p(x_{2})p(x_{4}).$ (IV.E.64)
Similarly, we obtain
$p(x_{1},x_{3},x_{4})$ $=p(x_{1})p(x_{3})p(x_{4}).$ (IV.E.65)
$p(x_{2},x_{3},x_{4})$ $=p(x_{2})p(x_{3})p(x_{4}).$ (IV.E.66)
Together with (IV.E.46)-(IV.E.48),
$p(x_{1})=p(x_{2})=p(x_{3}).$ (IV.E.67)
By (IV.E.49), (IV.E.50) and (IV.E.52), $X_{1}$, $X_{2}$ and $X_{3}$ are pairwise independent. Then by Lemma 1, we conclude that $X_{i}$ are uniformly distributed on $\mathcal{X}_{i}$ for $i\in N_{3}$ and $H(X_{1})=H(X_{2})=H(X_{3})=\log v$ where $v=|\mathcal{X}_{1}|=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|$.
Now we prove the “if ” part. Let $X_{2}$, $X_{3}$ and $X_{4}$ be mutually independent and distributed on $\mathbb{I}_{v}$, where $X_{2}$ and $X_{3}$ are uniform and $H(X_{4})=a$ for some $a\leq\log v$. Let $X_{1}=X_{2}+X_{3}+X_{4}\mod v$. Then for any $i\in\mathcal{X}_{1}$,
$p_{X_{1}}(i)=\sum_{x_{2},x_{3},x_{4}:\atop x_{2}+x_{3}+x_{4}\equiv i\mod v}p_{X_{2},X_{3},X_{4}}(x_{2}.x_{3},x_{4}).$ (IV.E.68)
Since $X_{2}$, $X_{3}$ and $X_{4}$ are mutually independent, we have
$p_{X_{1}}(i)=\sum_{x_{2},x_{3},x_{4}:\atop x_{2}+x_{3}+x_{4}\equiv i\mod v}p_{X_{2}}(x_{2})p_{X_{3}}(x_{3})p_{X_{4}}(x_{4})$ (IV.E.69)
Fix $x_{3}$ and $x_{4}$,
$p_{X_{1}}(i)=\sum_{j=0}^{v-1}\sum_{m=0}^{v-1}\sum_{x_{2}:\atop x_{2}+m+j\equiv i\mod v}p_{X_{2}}(x_{2})p_{X_{3}}(m)p_{X_{4}}(j).$ (IV.E.70)
Note that
$|\{x_{2}\in\mathcal{X}_{2}:x_{2}+m+j\equiv i\mod v\}|=1,$ (IV.E.71)
we have
$p_{X_{1}}(i)$ $=\sum_{j=0}^{v-1}\sum_{m=0}^{v-1}\dfrac{1}{v}p_{X_{3}}(m)p_{X_{4}}(j)$ (IV.E.72)
$=\dfrac{1}{v}\sum_{j=0}^{v-1}p_{X_{4}}(j)\sum_{m=0}^{v-1}p_{X_{3}}(m)$ (IV.E.73)
$=\dfrac{1}{v},$ (IV.E.74)
which implies $X_{1}$ is uniformly distributed on $\mathcal{X}_{1}$ and $H(X_{1})=\log v$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,b)$. ∎
Theorem IV-E.4.
For $F=(U_{3,4},U^{123}_{1,3})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if aa= $\log{v}$ for integer $v>0$ .
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{i4})$ $=H(X_{i})+H(X_{4}),i\in N_{3},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),k\neq 4,$
 $\qquad\qquad\text{ distinct }i,j,k\in N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.E.75)
$=p(x_{1},x_{2},x_{4})$ (IV.E.76)
$=p(x_{1},x_{3},x_{4})$ (IV.E.77)
$=p(x_{2},x_{3},x_{4}),$ (IV.E.78)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.E.79)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.E.80)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.E.81)
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (IV.E.82)
$p(x_{1},x_{2})p(x_{2},x_{3})$ $=p(x_{2})p(x_{1},x_{2},x_{3}),$ (IV.E.83)
$p(x_{1},x_{3})p(x_{2},x_{3})$ $=p(x_{3})p(x_{1},x_{2},x_{3}),$ (IV.E.84)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (IV.E.85)
$p(x_{1},x_{2})p(x_{2},x_{4})$ $=p(x_{2})p(x_{1},x_{2},x_{4}),$ (IV.E.86)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (IV.E.87)
$p(x_{1},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{1},x_{3},x_{4}),$ (IV.E.88)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.E.89)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}).$ (IV.E.90)
Since $p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})=p(x_{1},x_{3},x_{4})$, together with (IV.E.82), (IV.E.85) and (IV.E.87), we obtain
$p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4}).$ (IV.E.91)
By the same argument, we have
$p(x_{1},x_{2})$ $=p(x_{2},x_{3})=p(x_{2},x_{4}),$ (IV.E.92)
$p(x_{1},x_{3})$ $=p(x_{2},x_{3})=p(x_{3},x_{4}).$ (IV.E.93)
In light of (IV.E.91)-(IV.E.93),
$\quad\ p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4})$ (IV.E.94)
$=p(x_{2},x_{3})=p(x_{2},x_{4})=p(x_{3},x_{4}),$ (IV.E.95)
which implies the left hand side of (IV.E.82)-(IV.E.84) are equal. Then we conclude that $p(x_{1})=p(x_{2})=p(x_{3})$. By (IV.E.94), replacing $p(x_{1},x_{2})$ and $p(x_{1},x_{3})$ by $p(x_{1},x_{4})$ in (IV.E.82), we obtain
$p(x_{1},x_{4})p(x_{1},x_{4})=p(x_{1})p(x_{1},x_{2},x_{3}).$ (IV.E.96)
By (IV.E.79) and (IV.E.75), replacing $p(x_{1},x_{4})$ by $p(x_{1},x_{4})$, $p(x_{1},x_{2},x_{3})$ by $p(x_{1},x_{2},x_{3},x_{4})$, we have
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1})p^{2}(x_{4}).$ (IV.E.97)
By (IV.E.79), $X_{1}$ and $X_{4}$ are indenpendent. For any $x_{4}^{\prime}\in\mathcal{X}_{4}$ with $x_{4}^{\prime}\neq x_{4}$,
$p(x_{1},x_{4}^{\prime})=p(x_{1})p(x_{4}^{\prime})>0.$ (IV.E.98)
As
$p(x_{1},x_{4}^{\prime})=\sum_{x_{2},x_{3}}p(x_{1},x_{2},x_{3},x_{4}^{\prime}),$ (IV.E.99)
there exists $x_{2}^{\prime}\in\mathcal{X}_{2}$, $x_{3}^{\prime}\in\mathcal{X}_{3}$ such that $p(x_{1},x_{2}^{\prime},x_{3}^{\prime},x_{4}^{\prime})>0$. By the same argument, we have
$p(x_{1},x_{2}^{\prime},x_{3}^{\prime},x_{4}^{\prime})=p(x_{1})p^{2}(x_{4}^{\prime}).$ (IV.E.100)
Consider the tripartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}$ and $(x_{i},x_{j})\in E$ if and only if $p(x_{i},x_{j})>0$, ${i,j}\in N_{3},i\neq j$. Then, by Lemma 3, the probability mass of the triangles are equal in each connected component, which implies that $p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2}^{\prime},x_{3}^{\prime},x_{4}^{\prime})$. Equating (IV.E.97) and (IV.E.100), we have
$p(x_{4})=p(x_{4}^{\prime}).$ (IV.E.101)
It holds for any $x_{4}^{\prime}\in\mathcal{X}_{4}$, which implies that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}$ , and so $H(X_{4})=\log v$ where $v=|\mathcal{X}_{4}|$.
On the other hand, the “if” part of the theorem can be proved by Lemma 2 and the fact that $a=\log{v}$ on the ray $U_{3,4}$ and whole ray $U^{123}_{1,3}$ are entropic. ∎
Theorem IV-E.5.
For $F=(U_{3,4},U_{1,4})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a\geq 0,b>0$ or $(a,b)=(\log{v},0)$ for some positive integer $v$.
Figure 8: The face $(U_{3,4},U_{1,4})$
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),\text{ distinct }i,j,k\in N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.E.102)
$=p(x_{1},x_{2},x_{4})$ (IV.E.103)
$=p(x_{1},x_{3},x_{4})$ (IV.E.104)
$=p(x_{2},x_{3},x_{4}),$ (IV.E.105)
$p(x_{1},x_{2})p(x_{1},x_{3})$ $=p(x_{1})p(x_{1},x_{2},x_{3}),$ (IV.E.106)
$p(x_{1},x_{2})p(x_{2},x_{3})$ $=p(x_{2})p(x_{1},x_{2},x_{3}),$ (IV.E.107)
$p(x_{1},x_{3})p(x_{2},x_{3})$ $=p(x_{3})p(x_{1},x_{2},x_{3}),$ (IV.E.108)
$p(x_{1},x_{2})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{2},x_{4}),$ (IV.E.109)
$p(x_{1},x_{2})p(x_{2},x_{4})$ $=p(x_{2})p(x_{1},x_{2},x_{4}),$ (IV.E.110)
$p(x_{1},x_{4})p(x_{2},x_{4})$ $=p(x_{4})p(x_{1},x_{2},x_{4}),$ (IV.E.111)
$p(x_{1},x_{3})p(x_{1},x_{4})$ $=p(x_{1})p(x_{1},x_{3},x_{4}),$ (IV.E.112)
$p(x_{1},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{1},x_{3},x_{4}),$ (IV.E.113)
$p(x_{1},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{1},x_{3},x_{4}),$ (IV.E.114)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.E.115)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.E.116)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}).$ (IV.E.117)
Since $p(x_{1},x_{2},x_{3})=p(x_{1},x_{2},x_{4})=p(x_{1},x_{3},x_{4})$, by (IV.E.106),(IV.E.109) and (IV.E.112), we obtain
$p(x_{1},x_{2})p(x_{1},x_{3})=p(x_{1},x_{2})p(x_{1},x_{4})=p(x_{1},x_{3})p(x_{1},x_{4}),$ (IV.E.118)
which implies
$p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4}).$ (IV.E.119)
By the same argument, we have
$p(x_{1},x_{2})$ $=p(x_{2},x_{3})=p(x_{2},x_{4}).$ (IV.E.120)
$p(x_{1},x_{3})$ $=p(x_{2},x_{3})=p(x_{3},x_{4}).$ (IV.E.121)
Equating (IV.E.119)-(IV.E.121),
$\quad\ p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4})$ (IV.E.122)
$=p(x_{2},x_{3})=p(x_{2},x_{4})=p(x_{3},x_{4}),$ (IV.E.123)
which implies the left side of (IV.E.106)-(IV.E.117) are equal. Then with (IV.E.102)-(IV.E.105), we conclude that
$p(x_{1})=p(x_{2})=p(x_{3})=p(x_{4}).$ (IV.E.124)
Consider the tripartite graph $G=(V,E)$ with $V=\mathcal{X}_{1}\cup\mathcal{X}_{2}\cup\mathcal{X}_{3}$ and $(x_{i},x_{j})\in E$ if and only if $p(x_{i},x_{j})>0$, ${i,j}\in N_{3},i\neq j$. Assume $G$ has $t$ connected components. We denote the number of the vertices of $\mathcal{X}_{i}$ in the connected component $C_{j}$ by $n_{i}^{(j)},i=1,2,3,j=1,2,\cdots,t$ and the probability mass of $C_{j}$ by $p_{j}$. By Lemma 3, we obtain $n_{1}^{(j)}=n_{2}^{(j)}=n_{3}^{(j)}$ and so it can be simplified to $n^{(j)}$. The probability mass of the vertices, the edges and the triangles in $C_{j}$ are
$p(x_{1})=p(x_{2})=p(x_{3})=\dfrac{p_{j}}{n^{(j)}},$ (IV.E.125)
$p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{2},x_{3})=\dfrac{p_{j}}{(n^{(j)})^{2}},$ (IV.E.126)
$p(x_{1},x_{2},x_{3})=\dfrac{p_{j}}{(n^{(j)})^{3}},$ (IV.E.127)
respectively. Then
$H(X_{1})$ $=H(X_{2})=H(X_{3})$ (IV.E.128)
$=-\sum_{j=1}^{t}p_{j}\log\dfrac{p_{j}}{n^{(j)}}$ (IV.E.129)
$=H(p_{1},...,p_{t})+\sum_{j=1}^{t}p_{j}\log n^{(j)},$ (IV.E.130)
$H(X_{1},X_{2})$ $=H(X_{1},X_{3})=H(X_{2},X_{3})$ (IV.E.131)
$=-\sum_{j=1}^{t}p_{j}\log\dfrac{p_{j}}{(n^{(j)})^{2}}$ (IV.E.132)
$=H(p_{1},...,p_{t})+2\sum_{j=1}^{t}p_{j}\log n^{(j)},$ (IV.E.133)
$H(X_{1},X_{2},X_{3})=-\sum_{j=1}^{t}p_{j}\log\dfrac{p_{j}}{(n^{(j)})^{3}}$ (IV.E.134)
$=H(p_{1},...,p_{t})+3\sum_{j=1}^{t}p_{j}\log n^{(j)}.$ (IV.E.135)
As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=H(X_{2})=H(X_{3})$ (IV.E.136)
$=a+b,$ (IV.E.137)
$H(X_{1},X_{2})$ $=H(X_{1},X_{3})=H(X_{2},X_{3})$ (IV.E.138)
$=2a+b,$ (IV.E.139)
$H(X_{1},X_{2},X_{3})$ $=3a+b.$ (IV.E.140)
By (IV.E.128)-(IV.E.140), we have
$a$ $=\sum_{j=1}^{t}p_{j}\log n^{(j)},$ (IV.E.141)
$b$ $=H(p_{1},...,p_{t}).$ (IV.E.142)
By Lemma 2 and the fact that $a=\log v$ on the ray $U_{3,4}$, and the whole ray $U_{1,4}$ are entropic, all $\mathbf{h}=(a,b)\in F$ are entropic when $a=\log v$ for positive integer $v$, and $b\geq 0$. In the following, we only need to show that for all $\mathbf{h}=(a,b)\in F$ are entropic, when $a>0,b>0$ with $a\neq\log v$ for some integer $v>0$.
Consider $a\neq\log{v},b>0$. Let $c=\lfloor 2^{a}\rfloor$. Let $n^{(j)}=c$ for all $j\in N_{t-1}$, $n^{(t)}=c+l$ for some sufficiently large $l$ and $0<p_{i}<1$, $i=1,2,\dots$, $t$ be positive values satisfying $H(p_{1},...,p_{t})=b$ such that
$p_{t}=\dfrac{a-\log{\lfloor 2^{a}\rfloor}}{\log{(\lfloor 2^{a}\rfloor+l)}-\log{\lfloor 2^{a}\rfloor}}.$ (IV.E.143)
With such construction, we can check (IV.E.141) is stasified with
$\sum_{j=1}^{t}p_{j}\log n^{(j)}$ $=\sum_{j=1}^{t-1}p_{j}\log c+p_{t}\log{(c+l)}$ (IV.E.144)
$=(1-p_{t})\log c+p_{t}\log{(c+l)}$ (IV.E.145)
$=a.$ (IV.E.146)
Let $X^{(j)}_{1}$ $X^{(j)}_{2}$ and $X^{(j)}_{3}$ be mutually independent and uniformly distributed on $\mathbb{I}_{c}$ , and $X^{(j)}_{4}=X^{(j)}_{1}+X^{(j)}_{2}+X^{(j)}_{3}\mod c$ for $j\in N_{t-1}$. Let $X^{(t)}_{1}$, $X^{(t)}_{2}$ and $X^{(t)}_{3}$ be mutually independent and uniformly distributed on $\mathbb{I}_{c+l}$ and $X^{(t)}_{4}=X^{(t)}_{1}+X^{(t)}_{2}+X^{(t)}_{3}\mod(c+l)$. Assume $X^{(j)}_{i}$ is distributed on $\mathcal{X}^{(j)}_{i}$ for $i\in N_{4}$ and $j\in N_{t}$. Let $X_{i},i\in N_{4}$ be distributed on $\mathbb{I}_{tc+l}$ such that
$p_{X_{N_{4}}}(x_{i}+(j-1)c,i\in N_{4})$
$=$ $p_{j}p_{X^{(j)}_{1},X^{(j)}_{2},X^{(j)}_{3},X^{(j)}_{4}}(x_{1},x_{2},x_{3},x_{4})$ (IV.E.147)
for $j\in N_{t}$ and $(x_{1},x_{2},x_{3},x_{4})\in\mathcal{X}^{(j)}_{N_{4}}$. It can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $(a,b)$. Hence, all $\mathbf{h}=(a,b)\in F$ are entropic when $a>0,b>0$. The proof is accomplished. ∎
Theorem IV-E.5 can be considered as a corollary of [32, Lemma 4], which is restated as [36, Lemma 6].
IV-F Entropy functions on the faces involving $\hat{U}_{3,5}^{1}$
Theorem IV-F.1.
For $F=(\hat{U}_{3,5}^{1},U_{1,3}^{234})$ , $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log v$ for integer $v>0$.
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{1i})$ $=H(X_{1})+H(X_{i}),i\in\{2,3,4\},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{ijk}),\{i,j,k\}=\{2,3,4\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),\mid K\mid=2,$
$\qquad\qquad\qquad\qquad\{1\}\subseteq K\subseteq N_{4}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.F.1)
$=p(x_{1},x_{2},x_{4})$ (IV.F.2)
$=p(x_{1},x_{3},x_{4})$ (IV.F.3)
$=p(x_{2},x_{3},x_{4}),$ (IV.F.4)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.F.5)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.F.6)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.F.7)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.F.8)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.F.9)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (IV.F.10)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (IV.F.11)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (IV.F.12)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (IV.F.13)
Consider the tripartite graph $G=(V,E)$ with $V=\mathcal{X}_{2}\cup\mathcal{X}_{3}\cup\mathcal{X}_{4}$ and $(x_{i},x_{j})\in E$ if and only if $p(x_{i},x_{j})>0$, ${i,j}\in\{2,3,4\},i\neq j$. By (IV.F.8)-(IV.F.10) and Lemma 3, each connected component is a complete tripartite graph. By (IV.F.1), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (IV.F.11), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{2}).$ (IV.F.14)
Similarly, by (IV.F.1), (IV.F.2), (IV.F.12) and (IV.F.13), we obtain
$p(x_{1},x_{3},x_{4})=p(x_{1},x_{3}),$ (IV.F.15)
$p(x_{1},x_{3},x_{4})=p(x_{1},x_{4}).$ (IV.F.16)
In light of (IV.F.1)-(IV.F.7),
$p(x_{1234})=p(x_{1}.x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4})$ (IV.F.17)
$=$ $p(x_{1})p(x_{2})=p(x_{1})p(x_{3})=p(x_{1})p(x_{4}),$ (IV.F.18)
which implies $p(x_{2})=p(x_{3})=p(x_{4})$. By Lemma 3, the number of vertices in $\mathcal{X}_{i}$, $i=1,2,3$ are the same and the the probability mass of all of the vertices, the edges and the triangles are equal, respectively, in each connected component. Assume $G$ has $t$ connected components, $n^{(j)}$ and $p_{j}$ denote the number of the vertices and the probability mass of the connected component $C_{j}$, respectively. Then, the probability mass of the vertices, the edges and the triangles are $\frac{p_{j}}{n^{(j)}},\frac{p_{j}}{{(n^{(j)})}^{2}}$ and $\frac{p_{j}}{{(n^{(j)})}^{3}}$, respectively. Then,
$H(X_{2})$ $=H(X_{3})=H(X_{4})$ (IV.F.19)
$=-\sum_{i=1}^{t}n^{(j)}\dfrac{p_{j}}{n^{(j)}}\log{\dfrac{p_{j}}{n^{(j)}}}$ (IV.F.20)
$=H(p_{1},p_{2},...,p_{t})+\sum_{i=1}^{t}p_{j}\log{n^{(j)}},$ (IV.F.21)
$H(X_{2},X_{3})$ $=H(X_{2},X_{4})=H(X_{3},X_{4})$ (IV.F.22)
$=-\sum_{i=1}^{t}{(n^{(j)})}^{2}\dfrac{p_{j}}{{(n^{(j)})}^{2}}\log{\dfrac{p_{j}}{{(n^{(j)})}^{2}}}$ (IV.F.23)
$=H(p_{1},p_{2},...,p_{t})+2\sum_{i=1}^{t}p_{j}\log{n^{(j)}},$ (IV.F.24)
$H(X_{2}$ $,X_{3},X_{4})=-\sum_{i=1}^{t}{(n^{(j)})}^{3}\dfrac{p_{j}}{{(n^{(j)})}^{3}}\log{\dfrac{p_{j}}{{(n^{(j)})}^{3}}}$
$=H(p_{1},p_{2},...,p_{t})+3\sum_{i=1}^{t}p_{j}\log{n^{(j)}}.$ (IV.F.25)
As $\mathbf{h}\in F$, $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{2})$ $=H(X_{3})=H(X_{4})$ (IV.F.26)
$=a+b,$ (IV.F.27)
$H(X_{2},X_{3})$ $=H(X_{2},X_{4})=H(X_{3},X_{4})$ (IV.F.28)
$=2a+b,$ (IV.F.29)
$H(X_{2},X_{3},X_{4})$ $=3a+b.$ (IV.F.30)
By(IV.F.19)-(IV.F.30), we have
$a$ $=\sum_{j=1}^{t}p_{j}\log n^{(j)},$ (IV.F.31)
$b$ $=H(p_{1},...,p_{t}).$ (IV.F.32)
By (IV.F.4), $p(x_{2},x_{3},x_{4})=p(x_{1},x_{2},x_{3},x_{4})$, and so each triangle $(x_{2},x_{3},x_{4})$ can be colored by a unique $x_{1}\in\mathcal{X}_{1}$. For any $x_{2}\in\mathcal{X}_{2}$ in $C_{j}$ , there are ${(n^{(j)})}^{2}$ triangles containing the vertex $x_{2}$ based on the fact that each connected component is a complete tripartite graph. By (IV.F.17), $p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})$, which implies that these ${(n^{(j)})}^{2}$ triangles are colored differently. Note that $X_{1}$ and $X_{2}$ are independent by (IV.F.5), we have
$\mid\mathcal{X}_{1}\mid={(n^{(j)})}^{2}.$ (IV.F.33)
Note that it holds for all connected components of $G$, we obtain
$\mid\mathcal{X}_{1}\mid={(n^{(1)})}^{2}={(n^{(2)})}^{2}=...={(n^{(t)})}^{2},$ (IV.F.34)
which implies that
$n^{(1)}=n^{(2)}=\dots=n^{(t)}$ (IV.F.35)
and so they can be simplified to $v$. Then by (IV.F.31),
$a$ $=\sum_{j=1}^{t}p_{j}\log v=\log v\sum_{j=1}^{t}p_{j}=\log v.$ (IV.F.36)
Hence $a$ can just take the value of $\log v$, where $v$ is a positive integer.
As for “if” part of the theorem, it can be proved by Lemma 2 and the fact that $a=\log{v}$ on the ray $\hat{U}^{1}_{3,5}$ and whole ray $U_{1,3}^{234}$ are entropic. ∎
Theorem IV-F.2.
For $F=(\hat{U}_{3,5}^{1},U_{2,3}^{123})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a=\log{v_{1}},b=\log{v_{2}}$ for positive integer $v_{1},v_{2}$.
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),i<j,i,j\in N_{4},$
$H(X_{ik})+H(X_{jk})$ $=H(X_{k})+H(X_{i,j,k}),$
$\qquad\qquad\qquad\quad\{i,j,k\}=\{2,3,4\},$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),$
$\qquad\qquad\qquad\quad K=\{1,2\},\{1,3\}.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.F.37)
$=p(x_{1},x_{2},x_{4})$ (IV.F.38)
$=p(x_{1},x_{3},x_{4})$ (IV.F.39)
$=p(x_{2},x_{3},x_{4}),$ (IV.F.40)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.F.41)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.F.42)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.F.43)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (IV.F.44)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.F.45)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.F.46)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.F.47)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.F.48)
$p(x_{2},x_{4})p(x_{3},x_{4})$ $=p(x_{4})p(x_{2},x_{3},x_{4}),$ (IV.F.49)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (IV.F.50)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}).$ (IV.F.51)
By (IV.F.37)-(IV.F.40), (IV.F.50) and (IV.F.51), we have
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})=p(x_{1},x_{3}).$ (IV.F.52)
As $X_{1}$ is independent of $X_{2}$ and $X_{3}$,
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1})p(x_{2})=p(x_{1})p(x_{3}),$ (IV.F.53)
which implies $p(x_{2})=p(x_{3})$. Since $X_{2}$ and $X_{3}$ are independent by (IV.F.44), with Lemma 1, $X_{2}$ and $X_{3}$ are uniformly distributed on $\mathcal{X}_{2}$ and $\mathcal{X}_{3}$ and
$H(X_{2})=H(X_{3})=\log v,$ (IV.F.54)
where $v=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|$. Note that $X_{2}$ is independent of $X_{3}$ and $X_{4}$, replacing $p(x_{2},x_{3})$ and $p(x_{2},x_{4})$ by $p(x_{2})p(x_{3})$ and $p(x_{2})p(x_{4})$ in (IV.F.47), respectively, we have
$p(x_{2},x_{3},x_{4})=p(x_{2})p(x_{3})p(x_{4}).$ (IV.F.55)
Since $p(x_{1},x_{2},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$ by (IV.F.40), together with (IV.F.53) and (IV.F.55),
$p(x_{1})=p(x_{3})p(x_{4})=p(x_{2})p(x_{4}).$ (IV.F.56)
Note that $p(x_{2})=p(x_{3})=\frac{1}{v}$, we obtain
$p(x_{1})=\dfrac{1}{v}p(x_{4}).$ (IV.F.57)
Since $X_{1}$ and $X_{4}$ are independent, for any $x_{4}^{\prime}\in\mathcal{X}_{4}$ with $x_{4}^{\prime}\neq x_{4}$,
$p(x_{1},x_{4}^{\prime})=p(x_{1})p(x_{4}^{\prime}).$ (IV.F.58)
As $p(x_{1},x_{4}^{\prime})=\sum_{x_{2},x_{3}}p(x_{1},x_{2},x_{3},x_{4}^{\prime})$, there exists $x_{2}^{\prime}\in\mathcal{X}_{2},x_{3}^{\prime}\in\mathcal{X}_{3}$ such that $p(x_{1},x_{2}^{\prime},x_{3}^{\prime},x_{4}^{\prime})>0$. By the same argument, we have
$p(x_{1})=\dfrac{1}{v}p(x_{4}^{\prime}).$ (IV.F.59)
Equating (IV.F.57) and (IV.F.59),
$p(x_{4})=p(x_{4}^{\prime}),$ (IV.F.60)
which implies that $X_{4}$ is uniformly distributed on $\mathcal{X}_{4}$. Similarly, $X_{1}$ is uniformly distributed on $\mathcal{X}_{1}$. Assume $|X_{4}|=v_{1}$, we conclude that
$H(X_{1})=\log vv_{1},H(X_{4})=\log v_{1}.$ (IV.F.61)
For any $x_{1}\in\mathcal{X}_{1},x_{4}\in\mathcal{X}_{4}$, due to the independence of $X_{1}$ and $X_{4}$,
$p(x_{1},x_{4})=p(x_{1})p(x_{4})=\dfrac{1}{vv_{1}}\times\dfrac{1}{v_{1}}=\dfrac{1}{vv^{2}_{1}}.$ (IV.F.62)
As
$p(x_{1},x_{4})$ $=\sum_{x_{2},x_{3}}p(x_{1234})=\sum_{x_{2},x_{3}}p(x_{2},x_{3},x_{4})$ (IV.F.63)
$=\sum_{x_{2},x_{3}}p(x_{2})p(x_{3})p(x_{4})=\dfrac{k_{2}}{k^{2}k_{1}},$ (IV.F.64)
where $v_{2}=\mid\{(x_{2},x_{3})|\quad p(x_{1},x_{2},x_{3},x_{4})>0\}\mid$, we conclude that $v=v_{1}v_{2}$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{1})$ $=2a+b,$ (IV.F.65)
$H(X_{2})$ $=H(X_{3})=a+b,$ (IV.F.66)
$H(X_{4})$ $=a.$ (IV.F.67)
Together with (IV.F.54) and (IV.F.61), we obtain
$a=\log v_{1},b=\log\frac{v}{v_{1}}=\log v_{2},$ (IV.F.68)
which forms an outer bound on the entropy region on $F$, that is, $a$ and $b$ must take the value of $\log v_{1}$ and $\log v_{2}$, respectively, where $v_{1}$ and $v_{2}$ are positive integers.
The “if” part of the theorem can be immediately proved using Lemma 2, given the fact that $a=\log{v_{1}}$ in $\hat{U}^{1}_{3,5}$, and $b=\log{v_{2}}$ in $U_{2,3}^{123}$, both of which are entropic. ∎
Theorem IV-F.3.
For $F=(\hat{U}_{3,5}^{1},\mathcal{W}_{2}^{23})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if a+ba+b= $\log{v}$ for integer $v>0$.
Proof.
If $\mathbf{h}\in F$ is entropic, its characterizing random vector $(X_{i},i\in N_{4})$ satisfies the following information equalities,
$H(X_{N_{4}})$ $=H(X_{N_{4}-i}),\ i\in N_{4},$
$H(X_{ij})$ $=H(X_{i})+H(X_{j}),\{i,j\}\neq\{2,3\}$
$H(X_{23})+H(X_{k4})$ $=H(X_{k})+H(X_{2,3,4}),k\in\{2,3\}$
$H(X_{i\cup K})+H(X_{j\cup K})$ $=H(X_{K})+H(X_{ij\cup K}),\mid K\mid=2,$
$\hskip 108.12054pt\{1\}\subseteq K.$
For $(x_{i},i\in N_{4})\in\mathcal{X}_{N_{4}}$ with $p(x_{1234})>0$, above information equalities imply that the probability mass function satisfies
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (IV.F.69)
$=p(x_{1},x_{2},x_{4})$ (IV.F.70)
$=p(x_{1},x_{3},x_{4})$ (IV.F.71)
$=p(x_{2},x_{3},x_{4}),$ (IV.F.72)
$p(x_{1},x_{2})$ $=p(x_{1})p(x_{2}),$ (IV.F.73)
$p(x_{1},x_{3})$ $=p(x_{1})p(x_{3}),$ (IV.F.74)
$p(x_{1},x_{4})$ $=p(x_{1})p(x_{4}),$ (IV.F.75)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (IV.F.76)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (IV.F.77)
$p(x_{2},x_{3})p(x_{2},x_{4})$ $=p(x_{2})p(x_{2},x_{3},x_{4}),$ (IV.F.78)
$p(x_{2},x_{3})p(x_{3},x_{4})$ $=p(x_{3})p(x_{2},x_{3},x_{4}),$ (IV.F.79)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (IV.F.80)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (IV.F.81)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (IV.F.82)
By (IV.F.69)-(IV.F.72) and (IV.F.80)-(IV.F.82), we have
$p(x_{1234})=p(x_{1},x_{2})=p(x_{1},x_{3})=p(x_{1},x_{4}).$ (IV.F.83)
Since $X_{1}$ is independent of $X_{2},X_{3}$ and $X_{4}$ by (IV.F.73)-(IV.F.75), together with(IV.F.83)
$p(x_{1})p(x_{2})=p(x_{1})p(x_{3})=p(x_{1})p(x_{4}),$ (IV.F.84)
which implies $p(x_{2})=p(x_{3})=p(x_{4})$. Note that $X_{4}$ is independent of $X_{2}$ and $X_{3}$ by (IV.F.76) and (IV.F.77), by Lemma 1, $X_{i}$ is uniformly distributed on $\mathcal{X}_{i}$ and $H(X_{i})=\log v$ for $i=2,3,4$, where $v=|\mathcal{X}_{2}|=|\mathcal{X}_{3}|=|\mathcal{X}_{4}|$. As $\mathbf{h}\in F$ and $(X_{i},i\in N_{4})$ is its characterizing random vector, we have
$H(X_{2})=H(X_{3})=H(X_{4})=a+b.$ (IV.F.85)
Hence $a+b$ must take the value of $\log v$.
As for the “if” part of the theorem, we show that all the polymatroids $(a,b)$ satisfying $a+b=\log v$ are entropic. Let $p_{1},p_{2},\dots,p_{k}$ be positive value satisfying $H(p_{1},p_{2},\dots,p_{k})=a$. Let $X_{1}$ be distributed on $\mathbb{I}_{v^{2}}$ such that
$p_{X_{1}}(i)=\dfrac{p_{\lceil\frac{i+1}{k}\rceil}}{v},\text{ for }i=0,1,...,v^{2}-1.$ (IV.F.86)
Let $X_{2}$ be independent of $X_{1}$ and uniformly distributed on $\mathbb{I}_{v}$. Let
$X_{3}$ $=(\lfloor\frac{X_{1}}{v}\rfloor+X_{2})\mod v,$ (IV.F.87)
$X_{4}$ $=(X_{1}\mod v+X_{2})\mod v.$ (IV.F.88)
Then, it can be checked that the entropy function of such constructed $(X_{i},i\in N_{4})$ is $\mathbf{h}=(a,b)$. ∎
IV-G Non-entropic faces
Theorem IV-G.1.
For $F=(V^{12}_{8},E)$ with $E=U^{1}_{1,1},U^{3}_{1,1},U^{13}_{1,2},U^{134}_{1,3},U_{1,4},U^{123}_{2,3}$ or $U_{3,4}$, any $\mathbf{h}=(a,b)\in F$ is non-entropic if $a$ and $b$ are both positive.
Proof.
It can be immediately proven using [6, Thm. 11] that the extreme $1$-dimensional segment involving the top of the pyramid $(V^{\alpha}_{8})$ is not in $\overline{\Gamma_{4}^{*}}$. ∎